### v2

# Lissajous Varieties

Abstract
This paper studies affine algebraic varieties parametrized by sine and cosine functions, generalizing algebraic Lissajous figures in the plane. We show that, up to a combinatorial factor, the degree of these varieties equals the volume of a polytope. We deduce defining equations from rank constraints on a matrix with polynomial entries. We discuss applications of Lissajous varieties in dynamical systems, in particular the Kuramoto model. This leads us to study connections with convex optimization and Lissajous discriminants.
1 Introduction
A matrix $A\in\mathbb{Q}^{d\times n}$ defines a linear space ${\rm Row}(A)\subseteq\mathbb{C}^{n}$ by taking the $\mathbb{C}$-linear span of its rows. In this paper, we are interested in the algebraic varieties obtained by taking the coordinate-wise cosine of points in translates of ${\rm Row}(A)$. Concretely, let us define the map
${\rm cos}:\mathbb{C}^{n}\longrightarrow\mathbb{C}^{n},\quad(x_{1},\ldots,x_{n})\,\longmapsto\,(\cos(x_{1}),\ldots,\cos(x_{n})).$ (1)
For any vector $b\in\mathbb{C}^{n}$, we let ${\cal L}_{A,b}$ be the image of the affine-linear space $L_{A,b}={\rm Row}(A)-\frac{b\pi}{2}=\{x-\frac{b\pi}{2}\,:\,x\in{\rm Row}(A)\}\subseteq\mathbb{C}^{n}$ under the map $\rm cos$}. In symbols, we set
${\cal L}_{A,b}\,=\,{\rm cos}(L_{A,b}).$ (2)
Notice that ${\cal L}_{A,{\bf 0}}={\rm cos}({\rm Row}(A))$ and ${\cal L}_{A,{\bf 1}}={\rm sin}({\rm Row}(A))$, where ${\bf 0}\in\mathbb{C}^{n}$ and ${\bf 1}\in\mathbb{C}^{n}$ are the all-zeros and the all-ones vector respectively, and $\rm sin$} is the coordinate-wise sine map, analogous to (1). It is convenient to write ${\cal C}_{A}={\cal L}_{A,{\bf 0}}$ and ${\cal S}_{A}={\cal L}_{A,{\bf 1}}$ for these special cases.
The requirement that $A$ has rational entries ensures that ${\cal L}_{A,b}$ is an irreducible affine variety of dimension ${\rm rank}(A)$. In particular, the set $\cos(L_{A,b})\subseteq\mathbb{C}^{n}$ is Zariski closed in $\mathbb{C}^{n}$; see Lemma 2.1. Our first goal is to determine its degree and defining equations. We present some familiar examples and illustrate some features of ${\cal L}_{A,b}$.
Example 1.1.
The plane curves obtained from $A\in\mathbb{Q}^{1\times 2}$ and $b\in\mathbb{C}^{2}$ are known in the literature as Lissajous curves, which motivates the title of our paper. Such curves describe two objects driven in simple harmonic motion along the $x$- and $y$-axis [11]. In that context, one usually allows real entries for $A$, in which case ${\cal L}_{A,b}$ is not necessarily an algebraic curve. Moreover, one restricts to real vectors $b\in\mathbb{R}^{2}$ and focuses on the real points of ${\cal L}_{A,b}$.
Example 1.2.
The real points of $X=\{(x,y)\in\mathbb{C}^{2}\,:\,x^{2}+y^{2}=1\}$ form a circle of radius one, centered at the origin. The curve $X$ is parametrized by $(\cos(t),\sin(t))=(\cos(t),\cos(t-\frac{\pi}{2}))$. It is the Lissajous variety $X={\cal L}_{A,b}$ with $A=\begin{pmatrix}1&1\end{pmatrix}\in\mathbb{Q}^{1\times 2}$ and $b=(0,1)$.
Example 1.3.
Let $A=\left(\begin{smallmatrix}1&0&-1\\ -1&1&0\end{smallmatrix}\right)$. The surfaces ${\cal C}_{A}$ and ${\cal S}_{A}$ are parametrized as follows:
${\cal C}_{A}\,:\,(\cos(t_{1}-t_{2}),\,\cos(t_{2}),\,\cos(t_{1})),\quad{\cal S}_{A}\,:\,(\sin(t_{1}-t_{2}),\,\sin(t_{2}),\,-\sin(t_{1})).$
Figure 1 shows these surfaces.
Figure 1: The surfaces ${\cal C}_{A}$ (left) and ${\cal S}_{A}$ (right) from Example 1.3.
We compute that ${\cal C}_{A}$ is given by $1+2\,xyz-x^{2}-y^{2}-z^{2}=0$. This is Cayley’s cubic surface with four nodes. It is the algebraic boundary of the elliptope
${\cal E}_{3}\,=\,\left\{(x,y,z)\in\mathbb{R}^{3}\,:\,\begin{pmatrix}1&x&y\\ x&1&z\\ y&z&1\end{pmatrix}\text{ is positive semi-definite}\right\},$
a standard example of a spectrahedron in semi-definite programming [27]. The surface ${\cal S}_{A}$ is
$x^{4}+4\,x^{2}y^{2}z^{2}-2\,x^{2}y^{2}-2\,x^{2}z^{2}+y^{4}-2\,y^{2}z^{2}+z^{4}\,=\,0.$ (3)
It has degree six. As the parameter $b$ varies, the Lissajous varieties ${\cal L}_{A,b}$ form a family of surfaces. A generic fiber is reduced and irreducible of degree six. The special fiber for $b={\bf 0}$ is the cubic surface ${\cal C}_{A}$ with multiplicity two. This is a general phenomenon, see Section 2.
In the next few paragraphs we summarize our main contributions and sketch the outline of the paper. Since ${\cal L}_{A,b}$ only depends on the affine space $L_{A,b}$, one may replace $A$ by a matrix with the same row span without altering ${\cal L}_{A,b}$. Therefore, it is not restrictive to assume that $A$ is of rank $d$, and we may clear denominators so that $A$ has integer entries. Additionally, we make the following technical but equally non-restrictive assumption. The lattice $\mathbb{Z}^{d}$ is an abelian group with entry-wise addition. Let $\mathbb{Z}A\subseteq\mathbb{Z}^{d}$ be the subgroup generated by the columns $a_{1},\ldots,a_{n}\in\mathbb{Z}^{d}$ of $A$. We assume that $\mathbb{Z}A=\mathbb{Z}^{d}$. The following statement, proved and stated in more detail in Section 2, gives the degree of ${\cal L}_{A,b}$, i.e., the number of intersection points of ${\cal L}_{A,b}$ with a generic affine-linear space of complementary dimension $n-\dim({\cal L}_{A,b})$.
Theorem 1.4.
Let $A=\begin{pmatrix}a_{1}&a_{2}&\cdots&a_{n}\end{pmatrix}\in\mathbb{Z}^{d\times n}$ be such that ${\rm rank}(A)=d$ and $\mathbb{Z}A=\mathbb{Z}^{d}$. Let $P_{A}\subset\mathbb{R}^{d}$ be the polytope obtained as the convex hull of the lattice points $\{\pm a_{1},\ldots,\pm a_{n}\}\subset\mathbb{Z}^{d}$. For any $b\in\mathbb{C}^{n}$, the variety ${\cal L}_{A,b}$ is irreducible of dimension $d$. Moreover, for generic $b\in\mathbb{C}^{n}$, the degree of ${\cal L}_{A,b}$ is $\deg({\cal L}_{A,b})=2^{-{\rm CL}_{A}}\,d!\,{\rm vol}(P_{A})$, where ${\rm vol}(\cdot)$ denotes the euclidean volume and ${\rm CL}_{A}$ is the number of zero entries of a generic vector in the kernel of $A:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$.
The quantity $d!\,{\rm vol}(P_{A})$ is called the normalized volume of $P_{A}$. The polytope $P_{A}$ in Example 1.2 is the line segment [1,1][-1,1], with normalized volume two, which is the degree of the circle. The polygon $P_{A}$ in Example 1.3 is a hexagon with normalized volume $6=\deg{\cal S}_{A}$. The notion of “generic bb” in Theorem 1.4 will be made more precise in Section 2, and we give a formula for $\deg({\cal L}_{A,b})$ which holds for any $A,b$, but requires more notation (Theorem 2.5). We show that the hypersurface ${\cal C}_{A}$ obtained from the incidence matrix $A$ of the $n$-cycle graph $C_{n}$ is the cycle polynomial from [25, Section 4.2]. Using Theorem 2.5 and a result from [5], we prove a formula for the degree of the cycle polynomial (Proposition 2.10).
In Section 3, we construct a matrix with polynomial entries whose maximal minors form a set of set-theoretic defining equations of ${\cal L}_{A,b}$ (Theorem 3.1). In particular, a point $x^{*}\in\mathbb{C}^{n}$ belongs to ${\cal L}_{A,b}$ if and only if that matrix is not of full rank when evaluated at $x^{*}$.
Section 4 establishes the role of Lissajous varieties of type ${\cal S}_{A}$ in computing steady state angles for the Kuramoto equations of coupled oscillators. In this case, $A$ is the incidence matrix of a graph $G$ which encodes the coupling. We will see that the equilibrium angles correspond to the intersection points of ${\cal S}_{A}$ with an affine-linear space of the form $A\,x=\omega$. In particular, the degree of ${\cal S}_{A}$ bounds the number of isolated solutions.
In Section 5, we generalize the Kuramoto equations and construct dynamical systems whose steady state varieties are linear sections $A\,x=\omega$ of a Lissajous variety. Under certain assumptions on $\omega$, we show that one of the intersection points in ${\cal L}_{A,b}\cap\{A\,x=\omega\}$ corresponds to a stable equilibrium. It is the unique solution to a convex optimization problem. For varying $\omega$, these equilibria parametrize a subset of the Lissajous variety, called its positive part.
In Section 6, we study the discriminant of the equilibrium equations ${\cal L}_{A,b}\cap\{A\,x=\omega\}$ in the parameters $\omega$. We call this the Lissajous discriminant. We bound its degree and, when $A$ comes from a graph $G$, we describe its symmetries in terms of those of $G$. This is a first step in the bifurcation analysis of our dynamical system introduced in Section 5, see Example 6.7.
Related work.
Replacing $\rm cos$}” by $\rm exp$}” in (2), we obtain the affine toric variety $Y_{A}$ of the matrix $A$ up to scaling the coordinates by ${\rm exp}(b_{1}),\ldots,{\rm exp}(b_{n})$. Such scaled toric varieties appear in our study of the defining equations of ${\cal L}_{A,b}$, see Sections 3 and 4. The variety ${\cal C}_{A}={\cal L}_{A,{\bf 0}}$ is called a Chebyshev variety associated with $A$ in [1, Section 5]. This is motivated by the fact that for $d=1$, the curve ${\cal C}_{A}$ is alternatively obtained from a parametrization by Chebyshev polynomials of the first kind. Previous work on the algebraic geometry of Kuramoto equations includes [5, 12, 18]. These works use a different “algebraic geometrization” of the equations, see Equation (3) in [18] and Definition 1.4 in [12]. Some Lissajous discriminants were studied using machine learning techniques in [2, Section 5.3].
Acknowledgements.
We are grateful to Monique Laurent for helpful comments on a previous version of this manuscript. We thank Georgios Korpas, Hal Schenck and Rainer Sinn for useful conversations. This work has been supported by the European Union’s HORIZON–MSCA-2023-DN-JD programme under the Horizon Europe (HORIZON) Marie Skłodowska-Curie Actions, grant agreement 101120296 (TENORS).
2 Dimension and degree
As mentioned in the Introduction, if the matrices $A_{1}\in\mathbb{Q}^{d_{1}\times n},A_{2}\in\mathbb{Q}^{d_{2}\times n}$ have the same row span, then we have an equality of Lissajous varieties ${\cal L}_{A_{1},b}={\cal L}_{A_{2},b}$. In particular, after clearing denominators, we may assume that the matrix $A$ has integer entries. We fix $A\in\mathbb{Z}^{d\times n},b\in\mathbb{C}^{n}$.
Choosing coordinates on ${\rm Row}(A)$, we can parametrize ${\cal L}_{A,b}$ as follows:
$\phi_{A,b}\,:\,\mathbb{C}^{d}\longrightarrow\mathbb{C}^{n},\quad t=(t_{1},\ldots,t_{d})\longmapsto({\rm cos}(a_{1}\cdot t-b_{1}\tfrac{\pi}{2}),\ldots,{\rm cos}(a_{n}\cdot t-b_{n}\tfrac{\pi}{2})).$
Here $a_{j}\in\mathbb{Q}^{d}$ is the $j$-th column of $A$ and $a_{j}\cdot t$ is the standard dot product. We clearly have ${\cal L}_{A,b}={\rm im}\,\phi_{A,b}$. We obtain a rational parametrization of our Lissajous variety as follows. Set
$\beta_{\ell}\,=\,e^{-ib_{\ell}\tfrac{\pi}{2}},\,\,\ell=1,\ldots,n\quad\text{with}\quad i=\sqrt{-1}.$
Using Euler’s identity $\cos(\theta)=\tfrac{1}{2}(e^{i\theta}+e^{-i\theta})$, we see $\phi_{A,b}(t)=\psi_{A,b}(e^{it_{1}},\ldots,e^{it_{d}})$, with
$\psi_{A,b}\,:\,(\mathbb{C}^{*})^{d}\rightarrow\mathbb{C}^{n},\quad v=(v_{1},\ldots,v_{d})\longmapsto\Big(\frac{\beta_{1}v^{a_{1}}+\beta_{1}^{-1}v^{-a_{1}}}{2},\ldots,\frac{\beta_{n}v^{a_{n}}+\beta_{n}^{-1}v^{-a_{n}}}{2}\Big).$ (4)
Here $v^{a_{j}}=\prod_{k=1}^{d}v_{k}^{a_{kj}}$. We show below that ${\rm im}\,\phi_{A,b}={\rm im}\,\psi_{A,b}$, and therefore ${\cal L}_{A,b}={\rm im}\,\psi_{A,b}$.
Lemma 2.1.
The set ${\cal L}_{A,b}={\rm im}\,\psi_{A,b}={\rm im}\,\phi_{A,b}\subseteq\mathbb{C}^{n}$ is a closed affine subvariety of $\mathbb{C}^{n}$. Moreover, ${\cal L}_{A,b}$ is irreducible of dimension ${\rm rank}(A)$.
Proof 2.2.
The equality ${\rm im}\,\phi_{A,b}={\rm im}\,\psi_{A,b}$ follows from the fact that $\phi_{A,b}:\mathbb{C}^{d}\rightarrow\mathbb{C}^{n}$ is the composition of $\psi_{A,b}:(\mathbb{C}^{*})^{d}\rightarrow\mathbb{C}^{n}$ with the surjective map $\mathbb{C}^{d}\rightarrow(\mathbb{C}^{*})^{d}$ given by $t\mapsto(e^{it_{1}},\ldots,e^{it_{d}})$. The image ${\rm im}\,\phi_{A,b}={\rm im}\,\psi_{A,b}$ only depends on the row span of $A$. After applying integer row operations and column permutations to $A$, and after dropping rows whose entries are all zero, we obtain a matrix $\tilde{A}$ satisfying ${\rm im}\,\psi_{A,b}={\rm im}\,\psi_{\tilde{A},b}$ of the form
$\tilde{A}\,=\,\begin{pmatrix}a_{11}&0&\cdots&0&a_{1,r+1}&\cdots&a_{1,n}\\ 0&a_{22}&\cdots&0&a_{2,r+1}&\cdots&a_{2,n}\\ \vdots&\vdots&\ddots&\vdots&\vdots&\cdots&\vdots\\ 0&0&\cdots&a_{rr}&a_{r,r+1}&\cdots&a_{r,n}\end{pmatrix}\quad\in\,\mathbb{Z}^{r\times n},$
for some positive integers $a_{jj},j=1,\ldots,r$. Here $r\leq d$ is the rank of $A$. We want to show that the image of $\psi_{\tilde{A},b}\colon(\mathbb{C}^{*})^{r}\rightarrow\mathbb{C}^{n}$ is Zariski closed. Since $\mathbb{C}$ is algebraically closed, the Zariski closure of ${\rm im}\,\psi_{\tilde{A},b}$ equals its Euclidean closure. Hence, if $x^{*}\in\mathbb{C}^{n}$ lies in the closure of ${\rm im}\,\psi_{\tilde{A},b}$, there is a smooth path $\{x^{*}(t):t\in(0,1]\}$ contained in ${\rm im}\,\psi_{A,b}$ whose limit for $t\to 0$ is $x^{*}$. This lifts to a smooth path $v(t)=(v_{1}(t),\ldots,v_{r}(t))$ in $(\mathbb{C}^{*})^{r}$ with $\psi_{\tilde{A},b}(v(t))=x^{*}(t)$. We have
$x^{*}_{j}\,=\,\lim_{t\rightarrow 0}\tfrac{1}{2}(\beta_{j}\,v_{j}(t)^{a_{jj}}+\beta_{j}^{-1}\,v_{j}(t)^{-a_{jj}})\quad\text{ for }j=1,\ldots,r.$ (5)
Since the limit (5) lies in $\mathbb{C}$, we have $\lim_{t\rightarrow 0}v_{j}(t)\in\mathbb{C}^{*}$. Indeed, if the limit $\lim_{t\rightarrow 0}v_{j}(t)$ does not exist in $\mathbb{C}^{*}$, then $v_{j}(t)$ approaches 0 or $\infty$ for $t\rightarrow 0$. In both cases, $x_{j}(t)$ would not converge in $\mathbb{C}$. Hence, we have that $x^{*}=\lim_{t\rightarrow 0}\psi_{\tilde{A},b}(v(t))=\psi_{\tilde{A},b}(\lim_{t\rightarrow 0}v(t))\in{\rm im}\,\psi_{\tilde{A},b}$.
We have now shown that ${\cal L}_{A,b}={\rm im}\,\psi_{A,b}={\rm im}\,\psi_{\tilde{A},b}$ is closed in $\mathbb{C}^{n}$. Since ${\cal L}_{A,b}$ is unirational, it is irreducible. It is also clear that the projection of ${\rm im}\,\psi_{\tilde{A},b}$ to the first $r$ coordinates is surjective onto $\mathbb{C}^{r}$. Hence, we have $\dim{\cal L}_{A,b}=r={\rm rank}(A)$.
The equality ${\cal L}_{A,b}={\rm im}\,\psi_{A,b}$ implies an algorithm for computing the defining ideal of ${\cal L}_{A,b}$:
  1. 1.
    Work in the ring $\mathbb{C}[v_{1},\ldots,v_{d},w_{1},\ldots,w_{d},x_{1},\ldots,x_{n}]$ with $2d+n$ variables.
  2. 2.
    For each column $a_{i}$ of $A$, write $a_{i}=a_{i}^{+}-a_{i}^{-}$ with $a_{i}^{+},a_{i}^{-}\in\mathbb{N}^{d}$.
  3. 3.
    Define $I=\langle 2\,x_{j}-\beta_{1}v^{a_{j}^{+}}w^{a_{j}^{-}}-\beta_{1}^{-1}v^{a_{j}^{-}}w^{a_{j}^{+}},\,\,v_{k}w_{k}-1,\,\,j=1,\ldots,n,\,\,k=1,\ldots,d\rangle$.
  4. 4.
    Compute the elimination ideal $I\cap\mathbb{C}[x_{1},\ldots,x_{n}]$. The result equals $I({\cal L}_{A,b})$.
The ideal $I$ in step 3 has $n+d$ generators. Adding variables $w_{k}$ and imposing $v_{k}w_{k}-1$ is an effective way of working in the Laurent polynomial ring $\mathbb{C}[v_{1}^{\pm 1},\ldots,v_{d}^{\pm 1}]$: $w_{k}$ plays the role of $v_{k}^{-1}$. The correctness of this algorithm follows from the parametrization $\psi_{A,b}$. Notice that, for the ground field we can use any field extension of $\mathbb{Q}$ containing $\beta_{1},\ldots,\beta_{n}$. In particular, if $\beta_{\ell}$ is rational for $\ell=1,\ldots,n$, then one can replace $\mathbb{C}$ by $\mathbb{Q}$ in step 1. This happens when $b=\mathbf{0}$ and ${\cal L}_{A,b}={\cal C}_{A}$. If $b=\mathbf{1}$ and ${\cal L}_{A,b}={\cal S}_{A}$, then one can work over the field $\mathbb{Q}[i]=\mathbb{Q}[z]/\langle z^{2}+1\rangle$.
The integer matrix $A$ defines an affine toric variety $Y_{A}\subseteq\mathbb{C}^{n}$ obtained as the closure of the image of the Laurent monomial map $v\mapsto(v^{a_{1}},\ldots,v^{a_{n}})$, $v\in(\mathbb{C}^{*})^{d}$ [6, 26]. Its ideal is
$I_{A}\,=\,\langle y^{u}-y^{w}\,:\,u,w\in\mathbb{N}^{n}\,\,\text{and}\,\,A(u-w)=0\rangle\,\subseteq\,\mathbb{C}[y_{1},\ldots,y_{n}].$
A modified or scaled toric variety $Y_{A,\beta}$ is obtained from $A\in\mathbb{Z}^{d\times n}$, $\beta\in(\mathbb{C}^{*})^{n}$ as follows:
$Y_{A,\beta}\,=\,\{(\beta_{1}\,y_{1},\ldots,\beta_{n}\,y_{n})\,:\,y\in Y_{A}\}\,\subseteq\,\mathbb{C}^{n}.$
This variety is parametized by $v\mapsto(\beta_{1}v^{a_{1}},\ldots,\beta_{n}v^{a_{n}})$. One checks that if $I_{A}=I(Y_{A})=\langle y^{u_{1}}-y^{w_{1}},\ldots,y^{u_{r}}-y^{w_{r}}\rangle$, then $I_{A,\beta}=I(Y_{A,\beta})$ is generated by $\beta^{w_{k}}y^{u_{k}}-\beta^{u_{k}}y^{w_{k}},k=1,\ldots,r$.
Theorem 2.3.
The Lissajous variety ${\cal L}_{A,b}\subseteq\mathbb{C}^{n}$ is the image of the variety
${\cal Y}_{A,b}\,=\,\{(x,y)\in\mathbb{C}^{n}\times(\mathbb{C}^{*})^{n}\,:\,y\in Y_{A,\beta}\text{ and }x_{j}\,=\,\tfrac{1}{2}(y_{j}+y_{j}^{-1})\text{ for }j=1,\ldots,n\},$
under the coordinate projection $\pi_{A,b}:{\cal Y}_{A,b}\rightarrow\mathbb{C}^{n}$. Here $\beta=(e^{-ib_{1}\tfrac{\pi}{2}},\ldots,e^{-ib_{n}\tfrac{\pi}{2}})\in(\mathbb{C}^{*})^{n}$.
Proof 2.4.
The proof is an adaptation of the proof of [1, Theorem 5.2]. The map $y\mapsto(\tfrac{1}{2}(y_{1}+y_{1}^{-1}),\ldots,\tfrac{1}{2}(y_{n}+y_{n}^{-1}),y)$ induces an isomorphism $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}\simeq{\cal Y}_{A,b}$. Hence, we know from basic toric geometry that ${\cal Y}_{A,b}$ is a torus of dimension ${\rm rank}(A)$ [6, 26]. Composing this map with the coordinate projection $\pi_{A,b}$ and setting $y_{j}=\beta_{j}v^{a_{j}}$, we obtain precisely the map $\psi_{A,b}$ with image ${\cal L}_{A,b}$. This implies that ${\cal L}_{A,b}=\pi_{A,b}({\cal Y}_{A,b})$.
Deleting spurious rows if necessary, we may assume that $A$ has rank $d$. To state a degree formula for ${\cal L}_{A,b}$, let $\mathbb{Z}^{d}:\mathbb{Z}A]$ be the index of the lattice $\mathbb{Z}A$ generated by the columns of $A$ in the ambient lattice $\mathbb{Z}^{d}$. The degree $\deg\phi$ of a dominant morphism $\phi:X\rightarrow Y$ between irreducible varieties of dimension $d$ is the cardinality of a generic fiber.
Theorem 2.5.
Let $A\in\mathbb{Z}^{d\times n}$ be such that ${\rm rank}(A)=d$. Let $P_{A}\subset\mathbb{R}^{d}$ be the polytope obtained as the convex hull of the lattice points $\{\pm a_{1},\ldots,\pm a_{n}\}\subset\mathbb{Z}^{d}$. For any $b\in\mathbb{C}^{n}$, we have
$\deg({\cal L}_{A,b})\,=\,\frac{d!\,{\rm vol}(P_{A})}{\deg\pi_{A,b}\cdot[\mathbb{Z}^{d}:\mathbb{Z}A]}\,$,
where ${\rm vol}(\cdot)$ denotes the euclidean volume and $\pi_{A,b}:{\cal Y}_{A,b}\rightarrow{\cal L}_{A,b}$ is as in Theorem 2.3.
Proof 2.6.
This is a generalization of [1, Theorem 5.3]. The degree of ${\cal L}_{A,b}$ is the cardinality of
$S_{x}\,=\,\{x\in{\cal L}_{A,b}\,:\,c_{j0}+c_{j1}\,x_{1}+\cdots+c_{jn}\,x_{n}\,=\,0,\,\,\,j=1,\ldots,d\}$
for generic complex coefficients $c_{jk}$. We compare the set $S_{x}$ to the set $S_{v}$ given by
$S_{v}\,=\,\{v\in(\mathbb{C}^{*})^{d}\,:\,c_{j0}+\tfrac{c_{j1}}{2}(\beta_{1}v^{a_{1}}+\beta_{1}^{-1}v^{-a_{1}})+\cdots+\tfrac{c_{jn}}{2}(\beta_{n}v^{a_{n}}+\beta_{n}^{-1}v^{-a_{n}})\,=\,0,\,\,\,j=1,\ldots,d\}.$
By Lemma 2.1, we have $\psi_{A,b}(S_{v})=S_{x}$ for any choice of $c_{jk}$. Moreover, this correspondence is generically $\deg\psi_{A,b}$-to-one. The denominator $\deg\pi_{A,b}\cdot[\mathbb{Z}^{d}:\mathbb{Z}A]$ in our degree formula is the degree of $\psi_{A,b}$, since $\psi_{A,b}$ is the composition of the map $v\mapsto(\psi_{A,b}(v),\beta_{1}v^{a_{1}},\ldots,\beta_{n}v^{a_{n}})$ with $\pi_{A,b}$ (Theorem 2.3). Indeed, the first map has the same degree as $v\mapsto(\beta_{1}v^{a_{1}},\ldots,\beta_{n}v^{a_{n}})$, which parametrizes $Y_{A,\beta}$ and whose degree is $\mathbb{Z}^{d}:\mathbb{Z}A]$, see [26].
It remains to show that $S_{v}$ consists of $d!\,{\rm vol}(P_{A})$ points. This is an application of Kushnirenko’s theorem [13, Théorème III’], which predicts the maximal (and expected) number of solutions to a system of Laurent polynomial equations with identical monomial support. In order to apply this theorem, we must ensure that our equations are non-degenerate with respect to the polytope $P_{A}$. This means that none of the facial subsystems have a solution in $(\mathbb{C}^{*})^{d}$. By our assumption that ${\rm rank}(A)=d$, the centrally symmetric polytope $P_{A}$ is full-dimensional and it contains the origin in its interior. This implies that no face of $P_{A}$, except $P_{A}$ itself, contains both the lattice points $a_{j}$ and $a_{j}$. Therefore, there are no dependencies among the coefficients appearing in any facial subsystem. Hence, for generic $c_{jk}$, our equations are indeed non-degenerate and Kushnirenko’s upper bound $|S_{v}|\leq d!\,{\rm vol}(P_{A})$ is attained.
To complete the proof of Theorem 1.4, we must introduce some more notation. We write ${\rm Circ}(A)$ for the set of circuits of $A$. That is, ${\rm Circ}(A)$ is the set of all subsets of $n]=\{1,\ldots,n\}$ indexing a minimal set of linearly dependent columns of $A$. For any integer vector $m=(m_{1},\ldots,m_{n})\in\mathbb{Z}^{n}$, we define the support of $m$ as follows: ${\rm supp}(m)=\{j\in[n]\,:\,m_{j}\neq 0\}$. For each circuit $C\in{\rm Circ}(A)$ there is a unique (up to sign) integer vector $m^{C}=(m^{C}_{1},\ldots,m^{C}_{n})\in\mathbb{Z}^{n}$ of minimal length satisfying $A\cdot m_{C}=0$ and ${\rm supp}(m^{C})=C$. This vector encodes the unique linear relation between the columns indexed by the circuit $C$. A coloop of the matrix $A$ is an element $j\in[n]$ which does not belong to any circuit. We denote the set of coloops by ${\rm Coloops}(A)\subseteq[n]$, and its cardinality by ${\rm CL}_{A}=\#{\rm Coloops}(A)$. The integer ${\rm CL}_{A}$ is the number of zero entries of a generic vector in the kernel of $A$, as in Theorem 1.4.
Theorem 2.7.
Assume that $\mathbb{Z}A=\mathbb{Z}^{d}$, so that ${\rm rank}(A)=d$ and $\mathbb{Z}^{d}:\mathbb{Z}A]=1$. If for each circuit $C\in{\rm Circ}(A)$ the vector $b=(b_{1},\ldots,b_{n})\in\mathbb{C}^{n}$ satisfies
$m^{C}\cdot b\,=\,\sum_{j\in C}m^{C}_{j}\,b_{j}\quad\text{is not an even integer,}$ (6)
then $\deg\pi_{A,b}=2^{{\rm CL}_{A}}$ and the formula from Theorem 2.5 simplifies to $\deg({\cal L}_{A,b})=\frac{d!\,{\rm vol}(P_{A})}{2^{{\rm CL}_{A}}}$.
Proof 2.8.
To investigate the degree of $\pi_{A,b}$, pick a generic point $(x,y)\in{\cal Y}_{A,b}\subseteq\mathbb{C}^{n}\times(\mathbb{C}^{*})^{n}$ and notice that the only candidates for the points in the fiber $\pi_{A,b}^{-1}(\pi_{A,b}(x,y))$ are the $2^{n}$ points $(x,y_{1}^{\pm 1},\ldots,y_{n}^{\pm 1})$. For any subset $J\subseteq[n]$, let $\tilde{y}\in(\mathbb{C}^{*})^{n}$ be given by
$\tilde{y}_{j}\,=\,\begin{cases}y_{j}&j\notin J\\ y_{j}^{-1}&j\in J\end{cases}.$
To test whether $(x,\tilde{y})\in\pi_{A,b}^{-1}(\pi_{A,b}(x,y))$, we need to check whether $\tilde{y}\in Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$.
The variety $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$ is given by one binomial equation for each circuit of $A$:
$Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}\,=\,\{y\in(\mathbb{C}^{*})^{n}\,:\,y^{m^{C}}=\beta^{m^{C}}\,\,\,\text{for all}\,\,\,C\in{\rm Circ}(A)\}.$
If $J\subseteq{\rm Coloops}(A)$, then the coordinates $y_{j},j\in J$ do not appear in our binomial equations and $\tilde{y}$ lies on $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$. If $J\not\subseteq{\rm Coloops}(A)$ (in particular, $J\neq\emptyset$), pick a circuit $C$ such that $J\cap C\neq\emptyset$. If both $y$ and $\tilde{y}$ lie on $Y_{A,\beta}$, then
$y^{m^{C}}\cdot\tilde{y}^{m^{C}}\,=\,\prod_{j\in C\setminus J}y_{j}^{2\,m_{j}^{C}}\,=\,\beta^{\,2\,m^{C}}.$ (7)
The second equality in (7) holds for generic $y\in Y_{A,\beta}$ if and only if, for all $v\in(\mathbb{C}^{*})^{d}$, we have
$\prod_{j\in C\setminus J}(\beta_{j}v^{a_{j}})^{2\,m^{C}_{j}}\,=\,\beta^{\,2\,m^{C}},\quad\text{which implies}\quad v^{\,2\sum_{j\in C\setminus J}m^{C}_{j}\,a_{j}}\,=\,\prod_{j\in J\cap C}\beta_{j}^{\,2\,m^{C}_{j}}.$
In particular, we must have $\sum_{j\in C\setminus J}m_{j}^{C}\,a_{j}=0$. If $J\cap C\subsetneq C$, then this contradicts the fact that $C$ is a circuit. If $J\cap C=C$, then the equality fails if $\beta^{\,2\,m^{C}}\neq 1$. Taking the logarithm on both sides of this inequality, we obtain precisely the condition (6). We have shown that, under the condition (6) for every circuit $C\in{\rm Circ}(A)$, the points in $\pi_{A,b}^{-1}(\pi_{A,b}(x,y))$ are the points of the form $\tilde{y}$ corresponding to $J\subseteq{\rm Coloops}(A)$. Hence, the cardinality is $2^{{\rm CL}_{A}}$.
Remark 2.9.
When $b=\mathbf{0}$, the condition (6) is never satisfied. In fact, in this case, we have $\deg\pi_{A,b}\geq 2$, as the fiber $\pi_{A,\mathbf{0}}^{-1}(\pi_{A,\mathbf{0}}(x,y))$ contains $(x,y)$ and $(x,y^{-1})$ [1, Remark 5.4]. The second point is obtained as $\tilde{y}$ for $J=[n]$ in the notation of the proof of Theorem 2.7.
We end the section by applying our degree formula to a family of Chebyshev hypersurfaces which arises when studying elliptopes of cycle graphs [23, 25]. This is relevant in semidefinite completion problems, see for instance [15]. Rephrasing [25, Equation (28)] in our notation, the $n$-th cycle polynomial $\Gamma_{n}^{\prime}$ is the defining equation of the Chebyshev hypersurface ${\cal C}_{A_{n}}$, with
$A_{n}\,=\,\begin{pmatrix}1&0&\cdots&0&-1\\ 0&1&\cdots&0&-1\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&-1\end{pmatrix}\quad\in\,\,\mathbb{Z}^{(n-1)\times n}.$ (8)
Notice that $\Gamma_{n}^{\prime}$ is defined up to scale. Its degree was computed up to $n=11$ in [25, Table 1]. In the paragraph preceding Conjecture 4.9 in [25], it is stated that a closed formula for $\deg(\Gamma_{n}^{\prime})$ is not known. The next proposition provides such a formula.
Proposition 2.10.
The degree of the $n$-th cycle polynomial $\Gamma_{n}^{\prime}$ is given by
$\deg(\Gamma_{n}^{\prime})\,=\,\frac{n}{2}\binom{n-1}{\lfloor\tfrac{1}{2}(n-1)\rfloor}$. (9)
Proof 2.11.
By Theorem 2.5, we have $\deg(\Gamma_{n}^{\prime})=(\deg\pi_{A_{n},{\bf 0}})^{-1}(n-1)!\,{\rm vol}(P_{A_{n}})$. Let $y\in Y_{A_{n}}\cap(\mathbb{C}^{*})^{n}=\{y\in(\mathbb{C}^{*})^{n}\,:\,y_{1}\cdots y_{d+1}=1\}$ be a generic point on the affine toric hypersurface associated to $A_{n}$. The degree of $\pi_{A,{\bf 0}}$ is the number of points among $\{(y_{1}^{\pm 1},\ldots,y_{n}^{\pm 1})\}$ lying on $Y_{A_{n}}$. That number is two: only $(y_{1},\ldots,y_{n})$ and $(y_{1}^{-1},\ldots,y_{n}^{-1})$ lie on $Y_{A_{n}}$. The number $(n-1)!{\rm vol}(P_{A_{n}})$ was computed in [5, Theorem 14], and multiplying with $1/2$ gives (9).
3 Determinantal equations
Our goal in this section is to derive defining equations for ${\cal L}_{A,b}$ from rank conditions on a matrix with entries in $\mathbb{C}[x]$. Motivated by the equations $2\,x_{j}=y_{j}+y_{j}^{-1}$ appearing in Theorem 2.3, we work in the following setup. Let $K=\overline{\mathbb{C}(x)}$ be the algebraic closure of the field of rational functions in $x=(x_{1},\ldots,x_{n})$, and consider the following ideal:
$J\,=\,\langle 2\,x_{1}-y_{1}-y_{1}^{-1},\,\ldots,\,2\,x_{n}-y_{n}-y_{n}^{-1}\rangle\,\,\subset\,\,K[y^{\pm}]=K[y_{1}^{\pm 1},\ldots,y_{n}^{\pm 1}].$
The affine variety $V(J)$ defined by $J$ in $(K^{*})^{n}$ is zero-dimensional and consists of $2^{n}$ distinct points, each with multiplicity one. Hence, $J$ is radical and the quotient ${\cal A}=K[y^{\pm 1}]/J$ is a $K$-vector space of dimension $2^{n}$. Let $f_{j}=2\,x_{j}-y_{j}-y_{j}^{-1}$ be the $j$-th generator of $J$. We have
${\cal A}\,=\,\frac{K[y^{\pm 1}]}{\langle f_{1},\dotsc,f_{n}\rangle}\,\simeq\,\frac{K[y_{1}^{\pm 1}]}{\langle f_{1}\rangle}\otimes_{K}\,\cdots\,\otimes_{K}\frac{K[y_{n}^{\pm 1}]}{\langle f_{n}\rangle}\,.$ (10)
Below we write $[h]$ for the residue class of $h\in K[y^{\pm 1}]$ in $\cal A$}. Multiplication with an element $g\in K[y^{\pm 1}]$ gives a $K$-linear endomorphism $M_{g}:{\cal A}\rightarrow{\cal A}$, where $M_{g}([h])=[gh]$. Once we fix a $K$-basis for $\cal A$}, such a map is represented by a $2^{n}\times 2^{n}$ matrix with entries in $K$. We shall now describe such matrices for a basis compatible with the tensor product structure (10).
For a moment, set $n=1$. Let us fix the $K$-basis $\{[1],[y_{1}]\}$ for the algebra ${\cal A}=K[y_{1}^{\pm 1}]/\langle f_{1}\rangle$. We claim that, in this basis, multiplication with $y_{1}$ is given by the $2\times 2$ matrix
$M_{y_{1}}\,=\,\begin{pmatrix}0&-1\\ 1&2\,x_{1}\end{pmatrix}$.
Indeed, the first column reads $y_{1}\cdot 1]={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0}\cdot[1]+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1}\cdot[y_{1}]$, and the second column reads
$y_{1}\cdot y_{1}]\,=\,[y_{1}^{2}]\,=\,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1}\cdot[1]+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2\,x_{1}}\cdot[y_{1}],$
where the second equality follows from $y_{1}f_{1}]=0$. Multiplication with a general element $g\in K[y_{1}^{\pm 1}]$ is given by $M_{g}=g(M_{y_{1}})$, where $g(M_{y_{1}})$ denotes the matrix obtained by substituting $M_{y_{1}}$ for $y_{1}$ in the monomial expansion of $g$.
By (10), a general element $a\in{\cal A}$ is given by a finite sum $a=\sum_{k}a_{1}^{(k)}\otimes\dotsc\otimes a_{n}^{(k)}$, where
$a_{j}^{(k)}\in{\cal A}_{j}=K[y_{j}^{\pm 1}]/\langle f_{j}\rangle.$
Multiplication by a single variable, say $y_{1}$, satisfies
$M_{y_{1}}(a)\,=\,M_{y_{1}}\Big(\sum_{k}a_{1}^{(k)}\otimes\dotsc\otimes a_{n}^{(k)}\Big)\,=\,\sum_{k}M_{y_{1}}(a_{1}^{(k)})\otimes a_{2}^{(k)}\otimes\cdots\otimes a_{n}^{(k)}.$
Hence, fixing the basis $\{[1],[y_{j}]\}$ for ${\cal A}_{j}$ and the corresponding tensor product basis for ${\cal A}={\cal A}_{1}\otimes_{K}\cdots\otimes_{K}{\cal A}_{n}$, our previous observation gives
$M_{y_{1}}=\begin{pmatrix}0&-1\\ 1&2\,x_{1}\end{pmatrix}\otimes{\rm id}_{2}\otimes\dotsc\otimes{\rm id}_{2},$
where ${\rm id}_{2}$ is the $2\times 2$ identity matrix and $\otimes$ is the Kronecker product of matrices. The matrices $M_{y_{2}},\dotsc,M_{y_{n}}$ are found in an analogous way. By definition, our matrices $M_{y_{1}},\ldots,M_{y_{n}}$ are pairwise commuting. Hence, for any Laurent polynomial $g\in K[y^{\pm 1}]$, the matrix $g(M_{y_{1}},\ldots,M_{y_{n}})$ obtained by substituting $M_{y_{j}}$ for $y_{j}$ in the monomial expansion of $g$ is well-defined and represents the map $M_{g}$. Since the entries of both $M_{y_{j}}$ and $M_{y_{j}}^{-1}$ are polynomials in $x_{j}$, we have that $M_{g}\in\mathbb{C}[x]^{2^{n}\times 2^{n}}$ for any $g\in\mathbb{C}[y^{\pm 1}]\subset K[y^{\pm 1}]$. Here is the main theorem of this section.
Theorem 3.1.
Fix $A\in\mathbb{Z}^{d\times n}$ and $b\in\mathbb{C}^{n}$. Let $\beta\in(\mathbb{C}^{*})^{n}$ as usual, and $Y_{A,\beta}\subseteq\mathbb{C}^{n}$ be the corresponding scaled affine toric variety. If $g_{1},\dotsc,g_{r}$ generate the ideal of $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$ in $\mathbb{C}[y_{1}^{\pm 1},\ldots,y_{n}^{\pm 1}]$, then
${\cal L}_{A,b}\,=\,\{x\in\mathbb{C}^{n}:{\rm rank}(M_{g_{1}}|\dotsc|M_{g_{r}})<2^{n}\},$ (11)
where $M_{g_{k}}$ is a $2^{n}\times 2^{n}$-matrix with entries in $\mathbb{C}[x]$ representing multiplication by $g_{k}$ in $\cal A$}.
Remark 3.2.
It is easier to obtain generators for the ideal of $Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$ than for the toric ideal of $Y_{A,\beta}$. First, one computes a $\mathbb{Z}$-kernel of $A$. This can be found, for instance, via the Smith normal form of $A$. If this kernel is generated by $m_{1},\ldots,m_{r}\in\mathbb{Z}^{n}$, with $r=n-{\rm rank}(A)$, then the binomials $g_{1},\ldots,g_{r}$ in Theorem 3.1 are given by $y^{m_{k}}-\beta^{m_{k}},k=1,\ldots,r$ [26]. To avoid computing matrix inverses, one may clear denominators by writing $m_{k}=u_{k}-w_{k}$ with $u_{k},w_{k}\in\mathbb{N}^{n}$ and use $g_{k}=\beta^{w_{k}}y^{u_{k}}-\beta^{u_{k}}y^{w_{k}}$ instead.
Theorem 3.1 implies that a set of defining equations for ${\cal L}_{A,b}$ is given by the maximal minors of the $2^{n}\times(r\,2^{n})$-matrix $(M_{g_{1}}|\dotsc|M_{g_{r}})$ obtained by concatenating the matrices $M_{g_{j}}$ as block columns. In particular, if $A$ has rank n1n-1, then $Y_{A,b}$ is a hypersurface and its ideal $I_{A,b}=\langle g\rangle$ is principal. Hence, Theorem 3.1 provides a determinantal representation for ${\cal L}_{A,b}$. Before proving Theorem 3.1, we illustrate the statement in our running example.
Example 3.3.
Let $A$ be as in Example 1.3. Following the discussion above, we construct matrices representing multiplication with the variables in $K[y_{1}^{\pm 1},y_{2}^{\pm 1},y_{3}^{\pm 1}]/\langle f_{1},f_{2},f_{3}\rangle$:
$M_{y_{1}}=\begin{pmatrix}0&-1\\ 1&2\,x_{1}\end{pmatrix}\otimes{\rm id}_{2}\otimes{\rm id}_{2},\ \ M_{y_{2}}={\rm id}_{2}\otimes\begin{pmatrix}0&-1\\ 1&2\,x_{2}\end{pmatrix}\otimes{\rm id}_{2},\ \text{ and }\ M_{y_{3}}={\rm id}_{2}\otimes{\rm id}_{2}\otimes\begin{pmatrix}0&-1\\ 1&2\,x_{3}\end{pmatrix}$.
The toric ideal of $Y_{A,\beta}$ is generated by $g=y_{1}y_{2}y_{3}-\beta_{1}\beta_{2}\beta_{3}$. Evaluating this at our multiplication operators, we find $M_{g}=M_{y_{1}}M_{y_{2}}M_{y_{3}}-\beta_{1}\beta_{2}\beta_{3}\,{\rm id}_{8\times 8}$, which gives the following result:
$M_{g}\,=\,\begin{pmatrix}-\beta_{1}\beta_{2}\beta_{3}&0&0&0&0&0&0&-1\\ 0&-\beta_{1}\beta_{2}\beta_{3}&0&0&0&0&1&2x_{3}\\ 0&0&-\beta_{1}\beta_{2}\beta_{3}&0&0&1&0&2x_{2}\\ 0&0&0&-\beta_{1}\beta_{2}\beta_{3}&-1&-2x_{3}&-2x_{2}&-4x_{2}x_{3}\\ 0&0&0&1&-\beta_{1}\beta_{2}\beta_{3}&0&0&2x_{1}\\ 0&0&-1&-2x_{3}&0&-\beta_{1}\beta_{2}\beta_{3}&-2x_{1}&-4x_{1}x_{3}\\ 0&-1&0&-2x_{2}&0&-2x_{1}&-\beta_{1}\beta_{2}\beta_{3}&-4x_{1}x_{2}\\ 1&2x_{3}&2x_{2}&4x_{2}x_{3}&2x_{1}&4x_{1}x_{3}&4x_{1}x_{2}&8x_{1}x_{2}x_{3}-\beta_{1}\beta_{2}\beta_{3}\end{pmatrix}$
Setting $b=\mathbf{0}$ (i.e., $\beta=(1,1,1)$), the determinant evaluates to
$(\det M_{g})_{|\beta=(1,1,1)}\,=\,16\,(x_{1}^{2}-2x_{1}x_{2}x_{3}+x_{2}^{2}+x_{3}^{2}-1)^{2}.$
This is the square of the defining equation of Cayley’s cubic surface $\mathcal{C}_{A}={\cal L}_{A,\mathbf{0}}$. This computation shows that the equality (11) holds only set-theoretically; the maximal minors of $(M_{g_{1}}|\cdots|M_{g_{r}})$ do not necessarily generate a radical ideal. The equation (3) for $\mathcal{S}_{A}={\cal L}_{A,\mathbf{1}}$ is obtained by evaluating $\det M_{g}$ at $b=\mathbf{1}$ or $\beta=(-i,-i,-i)$. In this case, and for generic $\beta$, $\det M_{g}$ is an irreducible polynomial of degree six. The case $\beta=(1,1,1)$ is exceptional.
To prove Theorem 3.1, we first state a classical theorem, which describes the eigenvalues and left eigenvectors of a multiplication map. We refer to [10, Théorème 4.23] for details.
Theorem 3.4 (Eigenvalue, eigenvector theorem).
Let $\cal A$ be the coordinate ring of a zero-dimensional scheme with support $V=\{z_{1},\dotsc,z_{\delta}\}$. The eigenvalues of the multiplication map $M_{g}\colon\cal A\rightarrow\cal A$ are $g(z_{1}),\dotsc,g(z_{\delta})$, and the multiplicity of the eigenvalue $g(z_{j})$ equals the multiplicity of $z_{j}$. Moreover, for each $j=1,\ldots,\delta$, the evaluation map ${\rm ev}_{z_{j}}:{\cal A}\rightarrow\mathbb{C}$ given by ${\rm ev}_{z_{j}}([h])=h(z_{j})$ is a left eigenvector corresponding to the eigenvalue $g(z_{j})$.
Once a basis for the vector space $\cal A$} is fixed, the evaluation map ${\rm ev}_{z_{j}}:{\cal A}\rightarrow\mathbb{C}$ is represented by a row vector of length $\dim{\cal A}$. The eigenvalue relation is ${\rm ev}_{z_{j}}\,M_{g}\,=\,g(z_{j})\,{\rm ev}_{z_{j}}$.
Proof 3.5 (Proof of Theorem 3.1).
Let $W$ be the righthand side in (11). For $x^{*}\in\mathbb{C}^{n}$, let us write $J_{x^{*}}\subseteq\mathbb{C}[y^{\pm 1}]$ for the ideal generated by $(f_{1})_{x=x^{*}},\ldots,(f_{n})_{|x=x^{*}}$. The variety $V(J_{x^{*}})\subset(\mathbb{C}^{*})^{n}$ consists of at most $2^{n}$ points. Recall from Theorem 2.3 that ${\cal L}_{A,b}=\pi_{A,b}({\cal Y}_{A,b})$ and
$x^{*}\in\pi_{A,b}(\mathcal{Y}_{A,b})$ $\,\Longleftrightarrow\,\text{there is }z\in Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}\text{ such that }(f_{j})_{x=x^{*},y=z}=0\text{ for }j=1,\dotsc,n$
$\,\Longleftrightarrow\,\text{there is }z\in V(J_{x^{*}})\text{ such that }z\in Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$
$\,\Longleftrightarrow\,\text{there is }z\in V(J_{x^{*}})\text{ such that }g_{1}(z)=\cdots=g_{r}(z)=0.$
The ideal $J=\langle f_{1},\dotsc,f_{n}\rangle\subset K[y^{\pm 1}]$ is zero-dimensional and radical. By Theorem˜3.4, the eigenvalues of $M_{g_{j}}$ are given by $\{g_{j}(z):z\in V(J)\}\subset\overline{K}$. If there exists $z\in V(J_{x^{*}})$ such that $g_{1}(z)=\dotsc=g_{r}(z)=0$, then ${\rm ev}_{z}:\mathbb{C}[y^{\pm 1}]/J_{x^{*}}\rightarrow\mathbb{C}[y^{\pm 1}]/J_{x^{*}},[h]\mapsto h(z)$ is a common left eigenvector of $(M_{g_{1}})_{|x=x^{*}},\dotsc,(M_{g_{r}})_{|x=x^{*}}$ with eigenvalue zero. Hence, ${\rm ev}_{z}$ is represented by a row vector of length $2^{n}$, which is a left kernel vector of the concatenated matrix $(M_{g_{1}}|\cdots|M_{g_{r}})_{|x=x^{*}}$. This shows the inclusion $\pi_{A,b}({\cal Y}_{A,b})\subseteq W$, and hence ${\cal L}_{A,b}\subseteq W$.
For the reverse inclusion, suppose that ${\rm rank}(M_{g_{1}}|\dotsc|M_{g_{r}})_{|x=x^{*}}<2^{n}$ for some $x^{*}\in\mathbb{C}^{n}$. Then there exists a left null vector $v^{t}\in\mathbb{C}^{2^{n}}$. Applying column operations to the matrix $(M_{g_{1}}|\cdots|M_{g_{r}})_{|x=x^{*}}$, we may replace each $g_{j}$ by a random $\mathbb{C}$-linear combination $\tilde{g}_{j}$ of $g_{1},\ldots,g_{r}$. This has the effect that for each $z\in V(J_{x^{*}})$ and each $j=1,\ldots,r$, we have $\tilde{g}_{j}(z)=0$ if and only if $g_{1}(z)=\cdots=g_{r}(z)=0$, i.e., $\tilde{g}_{j}(z)=0\Leftrightarrow z\in Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}$. Since $v^{t}\cdot(M_{\tilde{g}_{1}}|\cdots|M_{\tilde{g}_{r}})_{|x=x^{*}}=0$, each of the matrices $M_{\tilde{g}_{j}}$ has a zero eigenvalue. Theorem 3.4 implies that for each $j$, $\tilde{g}_{j}(z^{(j)})=0$ for some $z^{(j)}\in V(J_{x^{*}})$. We conclude that $V(J_{x^{*}})\cap Y_{A,\beta}\cap(\mathbb{C}^{*})^{n}\neq\emptyset$ and $x^{*}\in\pi_{A,b}({\cal Y}_{A,b})$.
Proposition 3.6.
In the situation of Theorem 3.1, if $A$ has rank n1n-1 and $I_{A,\beta}=\langle g\rangle$, then
${\cal L}_{A,b}=\{x\in\mathbb{C}^{n}:\det(M_{g})=0\}.$ (12)
Moreover, if the prime ideal of ${\cal L}_{A,b}$ is $\langle f\rangle\subset\mathbb{C}[x]$, then $\det(M_{g})=c\,f^{\deg\pi_{A,b}}$ for some $c\in\mathbb{C}^{*}$.
Proof 3.7.
Equation (12) is an immediate consequence of Theorem 3.1. To compute the order of vanishing of $\det(M_{g})$ along ${\cal L}_{A,b}$, we introduce a small parameter $t$ and evaluate $M_{g}(x)$ at the point $x^{*}+t\cdot x_{0}$ for a generic point $x^{*}\in{\cal L}_{A,b}$ and a generic point $x_{0}\in\mathbb{C}^{n}$. By (12) we have
$\det M_{g}(x^{*}+t\cdot x_{0})\,=\,c\,f(x^{*}+t\cdot x_{0})^{k}\,=\,c_{1}\,t^{k}+O(t^{k+1})$ (13)
for some positive $k$ and $c_{1}\in\mathbb{C}^{*}$. By Theorem 3.4, the eigenvalues of $M_{g}(x^{*}+t\cdot x_{0})$ are the values $g(z(t))$ for $z(t)\in V(J_{x^{*}+t\cdot x_{0}})$. Therefore, we have $\det M_{g}(x^{*}+t\cdot x_{0})=\prod_{z(t)\in V(J_{x^{*}+t\cdot x_{0}})}g(z(t))$. By genericity of $x^{*}\in{\cal L}_{A,b}$, there are $\deg\pi_{A,b}$ eigenvalues for which $g(z(0))=0$. Hence, to conclude that $\det M_{g}(x^{*}+t\cdot x_{0})=c_{1}\,t^{\deg\pi_{A,b}}+O(t^{\deg\pi_{A,b}+1})$, it remains to show that for each eigenvalue with $g(z(0))=0$, we have $g(z(t))=\tilde{c}_{1}t+O(t^{2})$ for some $\tilde{c}_{1}\neq 0$. The coordinates of $z(t)$ satisfy $1-2(x_{j}^{*}+tx_{0j})z_{j}(t)+z_{j}(t)^{2}=0$. Solving this explicitly yields $z_{j}(t)=z_{j}(0)+t\,x_{0j}(1+x_{j}^{*}((x_{j}^{*})^{2}-1)^{-1/2})+O(t^{2})$. By genericity of $x_{0}$, the curve parametrized by $t\mapsto z(t)$ intersects $Y_{A,\beta}$ transversally at $t=0$. This implies that $g(z(t))$ has vanishing order one at $t=0$. We conclude that $k=\deg\pi_{A,b}$ in (13), as desired.
4 Kuramoto oscillators
The Kuramoto model is a system of ordinary differential equations, widely used to describe systems of coupled phase oscillators [14]. In this section, we explain how Lissajous varieties of type ${\cal S}_{A}$ show up in the study of its steady states. Let $G=(V,E)$ be a graph representing the coupling of a system of $m$ oscillators, where $m=|V|$. Let $n=|E|$ be the number of edges and, for $k=1,\ldots,m$, let $V_{k}$ be the set of vertices $v\in V$ adjacent to $v_{k}$. The Kuramoto model is a system of $m$ ordinary differential equations in $m$ unknown functions $\theta_{k}\colon{\mathbb{R}}\rightarrow\mathbb{R}:$
$\dot{\theta}_{k}\,=\,\omega_{k}+\sum_{j\in V_{k}}K_{kj}\,\sin{(\theta_{j}-\theta_{k})},\quad k\,=\,1,\ldots,m.$ (14)
The function $\theta_{k}(t)$ is the angle at vertex $k$ at time $t$, $\dot{\theta}_{k}$ is its derivative, $\omega_{k}\in\mathbb{R}$ is the natural frequency of the $k$-th oscillator, and $K_{kj}\in\mathbb{R}_{+}$ is the coupling strength of the edge $(k,j)\in E$.
Example 4.1.
Let $G=C_{3}$ be the triangle graph. Here $m=n=3$, and Equation˜14 reads:
$\dot{\theta}_{1}\,=\,K_{12}\sin{(\theta_{2}-\theta_{1})}+K_{13}\sin{(\theta_{3}-\theta_{1})}+\omega_{1},$
$\dot{\theta}_{2}\,=\,K_{12}\sin{(\theta_{1}-\theta_{2})}+K_{23}\sin{(\theta_{3}-\theta_{2})}+\omega_{2},$
$\dot{\theta}_{3}\,=\,K_{13}\sin{(\theta_{1}-\theta_{3})}+K_{23}\sin{(\theta_{2}-\theta_{3})}+\omega_{3}.$
The spring network associated with this graph is shown in Figure˜2. This is a mechanical illustration of the Kuramoto model [9]. The vertices of the graph are constrained to lie on a circle and are connected by spring-like edges. The Kuramoto model describes the angular velocity $\dot{\theta}_{k}$ of the particle at vertex $k$ as it moves around the circle. One sees from the equations that, for an equilibrium point to exist, we must have $\omega_{1}+\omega_{2}+\omega_{3}=0$.
[…]
Figure 2: The spring network of the triangle graph $C_{3}$.
We are interested in the steady states of Equation˜14. That is, we want to solve the trigonometric equations $\dot{\theta}_{k}=0$. For this we rephrase these equations as algebraic equations and use tools from computational algebraic geometry. This is much in the spirit of [5, 12, 18].
The approach taken in [5, 12, 18] and references therein is to substitute $x_{k}=\sin(\theta_{k})$ and $y_{j}=\cos(\theta_{j})$. The formula $\sin(\theta_{j}-\theta_{k})=\sin(\theta_{j})\cos(\theta_{k})-\cos(\theta_{j})\sin(\theta_{k})$ turns (14) into
$f_{k}\,=\,\omega_{k}+\sum_{v_{j}\in V_{k}}K_{kj}\,(x_{j}y_{k}-x_{k}y_{j}).$ (15)
Let $I_{\theta}=\langle x_{1}^{2}+y_{1}^{2}-1,\dotsc,x_{m}^{2}+y_{m}^{2}-1\rangle\subset\mathbb{C}[x,y]$ and $I_{G}=\langle f_{1},\dotsc,f_{m}\rangle\subset\mathbb{C}[x,y]$. In [12], the Kuramoto ideal is $I_{K}=I_{\theta}+I_{G}$ and the Kuramoto variety is its vanishing set $V(I_{K})\subseteq\mathbb{C}^{2m}$.
In this section we explore a different algebraic perspective. We assume that $G$ is a simple and connected graph and identify $V=[m]$. Let $A_{G}\in\mathbb{Z}^{m\times n}$ be the incidence matrix of $G$, defined as follows. If $k\to j$ is the $l$-th edge of $G$ with respect to an arbitrary fixed orientation and an arbitrary fixed ordering of the edges, then the $l$-th column of $A_{G}$ is $a_{l}=(e_{k}-e_{j})^{T}$, where $e_{k}$ is the $k$-th standard basis vector of $\mathbb{R}^{n}$. Since $G$ is connected, the rank of $A_{G}$ is m1m-1. The Lissajous variety ${\cal L}_{A_{G},\mathbf{1}}$ depends only on the row span of $A_{G}$. Therefore, unless otherwise specified, we let $A(G)$ be the matrix obtained by removing the last row of the incidence matrix $A_{G}$, and we set d=m1d=m-1. We refer to $A(G)$ as the reduced incidence matrix of $G$. Below, unless specified otherwise, we fix $G$ and write $A=A(G)\in\mathbb{Z}^{d\times n}$ for short. In particular, unless specified otherwise, ${\cal S}_{A}={\cal L}_{A(G),{\bf 1}}={\cal L}_{A_{G},{\bf 1}}$. We write $\omega\in\mathbb{R}^{d}$ for the vector of natural frequencies after dropping $\omega_{m}$. For simplicity, we shall assume in what follows that all constants $K_{ij}$ are equal to $K\in\mathbb{R}_{+}$. All statements are easily generalized to arbitrary $K_{ij}$.
With this setup, studying the steady states of Equation˜14 amounts to studying the intersection of the Lissajous variety ${\cal S}_{A}$ with an affine linear space. Indeed, after substituting $x_{l}=\sin(\theta_{k}-\theta_{j})$ in (14), the resulting equations are affine-linear. We obtain
$x\in{\cal S}_{A}\quad\text{and}\quad Ax=\omega/K.$ (16)
A steady state is recovered from $x$ satisfying (16) by computing the fiber $\phi_{A,\bf 1}^{-1}(x)$. A different approach is to set $v_{j}=e^{i\theta_{j}}$ and solve the following nonlinear equations in $(v_{1},\ldots,v_{d})$:
$\tfrac{\omega_{j}}{K}+\tfrac{a_{j1}}{2i}(v^{a_{1}}-v^{-a_{1}})+\cdots+\tfrac{a_{jn}}{2i}(v^{a_{n}}-v^{-a_{n}})\,=\,0,\,\,\,j=1,\ldots,d.$ (17)
The relation between these two approaches was exploited in the proof of Theorem 2.5, where the solution sets were denoted by $S_{x}$ and $S_{v}$ respectively, and we have $\psi_{A,\mathbf{1}}(S_{v})=S_{x}$.
Recall that the degree of ${\cal S}_{A}$ is given by the number of intersection points with a generic affine space of complementary dimension. The affine linear space $Ax=\omega/K$ has codimension $d$, which matches the dimension of ${\cal S}_{A}$. However, it is not generic, as both the variety and the affine space depend on the matrix $A$. On the other hand, if $Ax=\omega/K$ and ${\cal S}_{A}$ intersect in a finite number of isolated solutions, then this number is bounded above by the degree. The volume in Theorem 2.5 appeared as a bound on the number of isolated equilibria in [5].
Example 4.2.
Let $G=C_{3}$ be the triangle graph, as in Example 4.1. For cycle graphs, we choose the following edge ordering and orientation: $1\rightarrow 2,\,2\rightarrow 3,\,\ldots,\,n\rightarrow 1$. In the case $n=3$, the reduced incidence matrix $A$ with respect to this ordering coincides with $A$ from Example 1.3, hence yielding the same Lissajous variety ${\cal S}_{A}$. Since ${\rm rank}(A)$ is m1m-1, one of the equations in (14) can be dropped. For feasibility, we must have $\sum_{l=1}^{m}\omega_{l}=0$. We set the coupling strength $K$ to $1$, and choose $\omega=(\frac{1}{10},\frac{1}{5})$. We also fix $\theta_{m}=0$ and use $\theta_{1},\ldots,\theta_{m-1}=\theta_{d}$ as coordinates on ${\rm Row}(A)$. With these choices, Equation˜16 reads
$x^{4}+4x^{2}y^{2}z^{2}-2x^{2}y^{2}-2x^{2}z^{2}+y^{4}-2y^{2}z^{2}+z^{4}\,=\,10x-10z-1\,=\,-5x+5y-1\,=\,0.$
This system has six distinct real solutions, attaining the degree of ${\cal S}_{A}$. The $(x,y,z)$-coordinates of these solutions determine the sines of the steady state angles: $y=\sin(\theta_{2}),\,z=-\sin(\theta_{1})$. The cosines can be found by substituting these values appropriately in $f_{k}=0$, where $f_{k}$ is as in (15), and solve the resulting system of linear equations. Alternatively, one solves two Laurent polynomial equations in two unknowns given by (17), and finds $\theta_{j}=-i\log(v_{j})$.
Remark 4.3.
Let $G$ be the cycle graph $C_{n}$. With the choices of Example 4.2, its (reduced) incidence matrix has the same row span as the matrix $A_{n}$ from (8). The cycle polynomial $\Gamma_{n}^{\prime}$ is the defining equation of the Lissajous variety ${\cal C}_{A_{n}}={\cal L}_{A_{n},{\bf 0}}$.
We interpret the equations (16) from an optimization perspective. For this discussion, the matrix $A\in\mathbb{Z}^{d\times n}$ has rank $d$, and it does not necessarily come from a graph. We define a submanifold ${\cal S}_{A}^{+}\subset{\cal S}_{A}$ as follows. Consider the $d$-dimensional convex polytope $P={\rm Row}(A)\cap[-\tfrac{\pi}{2},\tfrac{\pi}{2}]^{n}$. obtained by intersecting a hypercube with the row span of $A$. We set
${\cal S}_{A}^{+}\,=\,\sin({\rm int}(P))\,\subset\,{\cal S}_{A}.$ (18)
That is, ${\cal S}_{A}^{+}$ is the image under $y\mapsto(\sin(y_{1}),\ldots,\sin(y_{n}))$ of the interior of $P\subset{\rm Row}(A)$. That map is an isomorphism of manifolds ${\rm int}(P)\simeq{\cal S}_{A}^{+}$ with inverse given by the coordinate-wise arcsine function ${\rm arcsin}\colon(-1,1)^{n}\rightarrow(-\tfrac{\pi}{2},\tfrac{\pi}{2})^{n},\,{\rm arcsin}(x_{1},\ldots,x_{n})=({\rm arcsin}(x_{1}),\ldots,{\rm arcsin}(x_{n}))$.
Theorem 4.4.
Let $A\in\mathbb{Z}^{d\times n}$ be of rank $d$. We have $x^{*}\in{\cal S}_{A}^{+}\cap\{AKx=\omega\}$ if and only if $x^{*}$ is the unique minimizer of the following convex optimization problem:
${\rm minimize}\,\sum_{j=1}^{n}\Big(x_{j}\,{\rm arcsin}(x_{j})+\sqrt{1-x_{j}^{2}}\Big),\quad{\rm subject\penalty 10000\ to}\quad Ax=\omega/K\,\,\text{and}\,\,x\in(-1,1)^{n}.$ (19)
Proof 4.5.
The function $g(t)=t\,{\rm arcsin}(t)+\sqrt{1-t^{2}}$ is strictly convex on the open interval (1,1)(-1,1). Hence, the objective function $\sum_{j=1}^{n}g(x_{j})$ is strictly convex on the feasible region of our optimization problem. If a minimizer $x^{*}\in(-1,1)^{n}$ exists, then it is unique, and it satisfies the first order optimality conditions
$\frac{\partial\,{\rm Lag}}{\partial x_{j}}\,=\,\frac{\partial}{\partial x_{j}}\Big(\sum_{j=1}^{n}g(x_{j})-\lambda^{t}(Ax-\omega/K)\Big)\,=\,0\quad\text{and}\quad Ax=\omega/K.$
Here $\rm Lag$} is the Lagrangian and $\lambda=(\lambda_{1},\ldots,\lambda_{d})$ are the Lagrange multipliers. The derivative of $g(t)$ is ${\rm arcsin}(t)$. Hence, the equations coming from partial derivatives of $\rm Lag$} with respect to $x_{j}$ are equivalent to ${\rm arcsin}(x)\in{\rm Row}(A)$, which implies ${\rm arcsin}(x)\in{\rm int}(P)$. Taking the coordinate-wise sine on both sides, we see that this is equivalent to $x\in{\cal S}_{A}^{+}$.
Motivated by Theorem 4.4, we introduce the following notation for the $A$-projection of ${\cal S}_{A}^{+}$:
$\Omega^{+}\,=\,\{A\,x\in\mathbb{R}^{d}\,:\,x\in{\cal S}_{A}^{+}\}\,=\,A({\cal S}_{A}^{+}).$
The optimization problem (19) has a unique minimizer if and only if $\omega/K\in\Omega^{+}$. The projection $A:{\cal S}_{A}^{+}\rightarrow\Omega^{+}$ is one-to-one, and the inverse is given by $\omega\mapsto{\cal S}_{A}^{+}\cap\{Ax=\omega\}$.
Example 4.6.
We set $K=1$ and use the matrix $A=\left(\begin{smallmatrix}1&0&-1\\ -1&1&0\end{smallmatrix}\right)$ associated with $G=C_{3}$. The manifold ${\cal S}_{A}^{+}$ is shown in blue in the left part Figure 3.
Figure 3: Positive regions for $x\in{\cal S}_{A}$ (left) and $\omega\in\mathbb{R}^{2}$ (right).
Its projection to $\mathbb{R}^{2}$ via $x\mapsto(x_{1}-x_{3},-x_{1}+x_{2})$ is $\Omega^{+}$, seen in Figure 3 (right). This set is contained in a hexagon $Q$, obtained as the $A$-projection of the cube $1,1]^{3}$. For $\omega\in{\rm int}(Q)$, the feasible region of (19) is an open line segment $\ell=\{Ax=\omega\}\cap(-1,1)^{n}$. For $\omega\in Q\setminus\Omega^{+}$, the objective function does not attain a minimum on $\ell$. For $\omega\in\Omega^{+}$, there is a unique minimizer given by $\ell\cap{\cal S}_{A}^{+}$.
We note that a different nonlinear (transcendental) optimization approach for $\omega=0$ is discussed in [16]. In Section 5, we shall establish a more general connection between Lissajous varieties and optimization. Before that, we conclude the present section with a remark on the stability of equilibria. Write the system of ODEs in Equation˜14 as $\dot{\theta}=\Phi(\theta)$. Let $J_{\Phi}(\theta)=\begin{pmatrix}\frac{\partial\Phi_{l}}{\partial\theta_{j}}(\theta)\end{pmatrix}_{lj}$ be the Jacobian matrix of $\Phi$. The matrix $J_{\Phi}(\theta)$ is symmetric, hence its eigenvalues are real. Because of the linear relation between the equations (14) pointed out above, $J_{\Phi}(\theta)$ is rank deficient – one of its eigenvalues is always zero. A solution $\theta^{*}$ to $\Phi(\theta)=0$ is called linearly stable if all other eigenvalues of $J_{\Phi}(\theta^{*})$ are negative. If $\omega/K\in\Omega^{+}$, then the equilibrium $\theta^{*}$ corresponding to the minimizer $x^{*}$ from Theorem 4.4 is linearly stable.
Proposition 4.7.
Let $A$ be the reduced incidence matrix of $G$ and let $\omega/K\in\Omega^{+}$. Then, the unique minimizer described in Theorem˜4.4 yields a linearly stable equilibrium $\theta^{*}$ of the corresponding Kuramoto model. This is the unique vector $\theta^{*}$ satisfying $A^{t}\theta^{*}={\rm arcsin}(x^{*})$.
Proof 4.8.
If $\omega/K\in\Omega^{+}$, then $A^{t}\theta^{*}\in{\rm int}(P)\subset{\rm Row}(A)$. In particular, $\cos(a_{l}\cdot\theta^{*})>0$ for all $l=1,\dotsc,n$. The statement then follows from [12, Lemma 3.2].
While solving the equations (17) leads to all potential equilibria, there may be many “spurious” solutions which have no physical meaning. Proposition 4.7 says that if $\omega/K\in\Omega^{+}$, then a linearly stable solution can be found more efficiently via convex optimization. We now consider a larger graph to illustrate the computational range of this technique.
Example 4.9.
The graph shown in Figure 4 appears in [12, Example 3.1]. The corresponding Lissajous variety ${\cal S}_{A}$ has degree 832. For $\omega=(0.1,\ldots,0.1)$, the unique minimizer in (19) is
$x^{*}\,=\,(-0.050,0.150,0.083,0.017,-0.050,0.150,0.017,0.083,-0.067,0.200,0.067,0.200),$
here rounded to three decimal places. This can be computed using, for instance, JuMP.jl [17]. The optimum is found in less than 0.01 seconds. Solving $A^{t}\theta^{*}={\rm arcsin}(x^{*})$ we find that
$\theta^{*}=(0.151,0.285,0.151,0.285,0.201,0.268,0.201),$
which is a linearly stable equilibrium of the Kuramoto equations. We verify that $x\in{\cal S}_{A}^{+}\cap\{A\,x=\omega\}$ and $\omega\in\Omega^{+}$ by checking that $\omega=A\,{\rm sin}(A^{t}\theta^{*})$, and $A^{t}\theta^{*}\in[-\tfrac{\pi}{2},\tfrac{\pi}{2}]^{12}$.
Figure 4: A graph and its reduced incidence matrix.
5 Positive points and convex optimization
In Theorem 4.4 we showed that the open subset ${\cal S}_{A}^{+}\subset{\cal S}_{A}={\cal L}_{A,\bf 1}$ parametrizes all solutions of the optimization problem (19) for $\omega\in\Omega^{+}$. In this section we prove a similar characterization for an arbitrary Lissajous variety ${\cal L}_{A,b}$. In particular, we generalize Theorem 4.4 and Proposition 4.7. The Kuramoto model is generalized by the dynamical system
$\dot{\theta}\,=\,A\,\phi_{A,b}(\theta)-\omega,$ (20)
where $A\in\mathbb{Q}^{d\times n}$ has rank $d$, $b\in\mathbb{R}^{n}$, and $\omega\in\mathbb{R}^{d}$. The map $\phi_{A,b}:[-\pi,\pi]^{d}\rightarrow[1,1]^{n}$ is
$\phi_{A,b}(\theta)=(\cos(a_{1}\cdot\theta-b_{1}\,\tfrac{\pi}{2}),\ldots,\cos(a_{1}\cdot\theta-b_{n}\,\tfrac{\pi}{2})).$
Investigating the equilibria of this dynamical system leads us to solve $A\,\phi_{A,b}(\theta)-\omega=0$. We complexify $\phi_{A,b}:\mathbb{C}^{d}\rightarrow\mathbb{C}^{n}$ and refer to the solutions $\theta\in\mathbb{C}^{d}$ as steady states or equilibria. The steady states correspond to points in the intersection ${\cal L}_{A,b}\cap\{Ax=\omega\}$. We shall define ${\cal L}_{A,b}^{+}\subset{\cal L}_{A,b}\cap(-1,1)^{n}$ such that there is at most one intersection point in ${\cal L}_{A,b}^{+}\cap\{Ax=\omega\}$. Moreover, if such a point exists, then it is the solution to a convex optimization problem.
The affine linear space $L_{A,b}$ is defined as $L_{A,b}={\rm Row}(A)-\tfrac{b\pi}{2}$. Its image under the coordinate-wise cosine is ${\cal L}_{A,b}$. The appropriate generalizations of ${\cal S}_{A}^{+}$ and $\Omega^{+}$ are as follows:
${\cal L}_{A,b}^{+}\,=\,\cos(L_{A,b}\cap(0,\pi)^{n})\,\subset\,{\cal L}_{A,b},\quad\quad\Omega^{+}_{A,b}\,=\,\{A\,x\in\mathbb{R}^{d}\,:\,x\in{\cal L}_{A,b}^{+}\}\,=\,A({\cal L}_{A,b}^{+}).$
One checks that this is consistent with the previous section, in that ${\cal L}_{A,\bf 1}^{+}={\cal S}_{A}^{+}$ and $\Omega_{A,{\bf 1}}^{+}=\Omega^{+}$.
Restricting to the real points of $L_{A,b}$, we have that ${\rm cos}(L_{A,b}\cap\mathbb{R}^{n})=\cos(L_{A,b}\cap[-\pi,\pi]^{n})$. This is a subset of ${\cal L}_{A,b}(\mathbb{R})$. The ‘++’ in our notation is motivated by the fact that ${\cal L}_{A,b}^{+}$ is the image under the cosine map of all “positive” tuples of angles in $L_{A,b}\cap(0,\pi)^{n}$.
Theorem 5.1.
Let $A\in\mathbb{Q}^{d\times n}$ be of rank $d$. We have $x^{*}\in{\cal L}_{A,b}^{+}\cap\{Ax=\omega\}$ if and only if $x^{*}$ is the unique minimizer of the following convex optimization problem:
${\rm minimize}\,\,\frac{\text{-}\pi}{2}b^{t}x-\sum_{j=1}^{n}\Big(x_{j}\,{\rm arccos}(x_{j})-\sqrt{1-x_{j}^{2}}\Big),\quad{\rm s.\,t.}\quad Ax=\omega\,\,\text{and}\,\,x\in(-1,1)^{n}.$ (21)
In particular, a minimizer exists if and only if $\omega\in\Omega_{A,b}^{+}$, and in that case it is unique.
Proof 5.2.
The first and second order derivatives of $g(t)=-t\,{\rm arccos}(t)+\sqrt{1-t^{2}}$ are
$g^{\prime}(t)\,=\,-{\rm arccos}(t)\quad\text{and}\quad g^{\prime\prime}(t)=(1-t^{2})^{-1/2}.$
Hence, the objective function in (21) is strictly convex on $(-1,1)^{n}$, and so is its restriction to the feasible region. The rest of the proof is analogous to that of Theorem 4.4. The method of Lagrange multipliers gives the first order optimality conditions ${\rm arccos}(x)\in{\rm Row}(A)-\tfrac{\pi b}{2}$ and $Ax=\omega$, which is equivalent to $x\in{\cal L}_{A,b}^{+}\cap\{Ax=\omega\}$.
Remark 5.3.
Theorem 5.1 is inspired by analogous convex optimization problems in which the constraint $x\in(-1,1)^{n}$ is replaced by $x\in\mathbb{R}^{n}_{+}$ and the objective function is a strictly convex function $G(x)$ on the positive orthant. Often $G$ is of the form $G(x)=\sum_{j=1}^{n}g(x_{j})$. Appropriate choices of $g$ lead naturally to semi-algebraic descriptions of the unique minimizer, similar to $x^{*}={\cal L}_{A,b}\cap\{Ax=\omega\}$ in Theorem 5.1. For $g(t)=\log(t)$, the Lissajous variety is replaced by a reciprocal linear space [7]. The function $g(t)=t\log(t)-t$ naturally leads to positive toric varieties [24]. If the universal barrier function $G(x)$ of the feasible polytope is minimized instead, then one intersects $\{Ax=\omega\}$ with the Santaló patchwork [21].
Remark 5.4.
The coordinates of the minimizer $x^{*}$ of (21) are algebraic functions of $\omega$. Their minimal polynomial in $\mathbb{Q}(\omega)[x_{j}]$ has degree at most $\deg({\cal L}_{A,b})$, see the formula stated in Theorem 2.5. This follows from the fact that $x^{*}\in{\cal L}_{A,b}\cap\{Ax=\omega\}$ by Theorem 5.1.
Write the equations (20) as $\dot{\theta}=\Phi_{A,b}(\theta)$ and let $J_{\Phi_{A,b}}(\theta)$ be the $d\times d$ Jacobian matrix. We say that a steady state solution $\theta^{*}$ is linearly stable if all eigenvalues of $J_{\Phi}(\theta^{*})$ are negative.
Proposition 5.5.
Let $\omega\in\Omega_{A,b}^{+}$ and let $x^{*}\in{\cal L}_{A,b}^{+}$ be the unique minimizer of the optimization problem (21). The unique solution $\theta^{*}$ of the linear equations $A^{t}\theta^{*}={\rm arccos}(x^{*})+\tfrac{\pi b}{2}$ is a linearly stable steady state solution of the dynamical system (20).
Proof 5.6.
Note that $\theta^{*}$ is a steady state solution by construction: $A\phi_{A,b}(\theta^{*})=A\,x^{*}=\omega$. The Jacobian matrix $J_{\Phi_{A,b}}$ has the following explicit expression:
$J_{\Phi_{A,b}}(\theta)\,=\,-A\,\,{\rm diag}(\sin(a_{1}\cdot\theta-b_{1}\tfrac{\pi}{2}),\ldots,\sin(a_{n}\cdot\theta-b_{n}\tfrac{\pi}{2}))\,\,A^{t}.$
Since ${\rm arccos}(x^{*})=A^{t}\theta^{*}-\frac{\pi b}{2}\in(0,\pi)^{n}$, the diagonal matrix in this expression has positive diagonal entries for $\theta=\theta^{*}$. Hence, $J_{\Phi_{A,b}}(\theta^{*})$ is negative definite.
Example 5.7.
The Lissajous curve ${\cal L}_{A,b}$ for the data $A=\begin{pmatrix}1&1\end{pmatrix}$, $b=(0,1)$ is the circle with defining equation $x^{2}+y^{2}=1$, see Example 1.2. The semi-algebraic set ${\cal L}_{A,b}^{+}$ is the segment of the curve contained in the (,+)(-,+) quadrant, see Figure 5.
Figure 5: The line $x+y=\omega$ has precisely one intersection point with ${\cal L}_{A,b}^{+}$ for $\omega\in(-1,1)$.
The projection $\Omega_{A,b}^{+}=A({\cal L}_{A,b}^{+})$ is the open line segment (1,1)(-1,1). For $\omega\in\Omega_{A,b}^{+}$, the line $x+y=\omega$ has one intersection point with ${\cal L}_{A,b}^{+}$. This is the unique solution to (21). The differential equation
$\dot{\theta}\,=\,\cos(\theta)+\sin(\theta)-\omega,\quad\omega\in\Omega_{A,b}^{+}$
has two steady state solutions $\theta^{*}$. Only one of them is stable. For concreteness, set $\omega=0.6$. The stable equilibrium is $\theta^{*}\approx 1.918$. Its image $(\cos(\theta^{*}),\sin(\theta^{*}))$ is ${\cal L}_{A,b}^{+}\cap\{x+y=0.6\}$.
6 Lissajous discriminants
Varying the natural frequencies $\omega$ affects the number of real solutions to the systems of polynomial equations (16) and (17). In the context of dynamical systems, varying $\omega$ influences the nature of the equilibria, which is the topic of bifurcation analysis. Investigating this leads us to study a discriminant locus in the $\omega$ parameters. More precisely, we shall define a variety $\nabla_{A,b}\subset\mathbb{C}^{d}$, which is expected to be a hypersurface, such that the number of real solutions to
$\tfrac{a_{j1}}{2}(\beta_{1}v^{a_{1}}+\beta_{1}^{-1}v^{-a_{1}})+\cdots+\tfrac{a_{jn}}{2}(\beta_{n}v^{a_{n}}+\beta_{n}^{-1}v^{-a_{n}})-\omega_{j}\,=\,0,\quad j=1,\ldots,d$ (22)
is constant for $\omega$ in each connected component of $\mathbb{R}^{d}\setminus\nabla_{A,b}$. For $b=\mathbf{1}$ ($\beta=-i\cdot\mathbf{1}$), these are the equations (17). Taking cues from standard discriminant analysis, $\nabla_{A,b}$ should consist of points $\omega\in\mathbb{C}^{d}$ for which two complex solutions of (22) collide. We now make this precise.
We continue to assume that the matrix $A\in\mathbb{Z}^{d\times n}$ has full rank $d$. Note that we can write the equations (22) in a compact way as follows: $A\,\psi_{A,b}(v)=\omega$, where $\psi_{A,b}$ is as in (4). Consider the incidence variety $W_{A,b}=\{(v,\omega)\in(\mathbb{C}^{*})^{d}\times\mathbb{C}^{d}\,:\,A\,\psi_{A,b}(v)=\omega\}$. The fiber of
${\rm pr}_{\omega}\,:\,W_{A,b}\,\longrightarrow\,\mathbb{C}^{d},\quad(v,\omega)\longmapsto\omega$ (23)
over $\omega\in\mathbb{C}^{d}$ consists of the solutions $v$ to (22) for that fixed value of $\omega$. The toric Jacobi matrix of $v\mapsto A\,\psi_{A,b}(v)$ is the $d\times d$ matrix given by
$J_{A,b}(v)\,=\,\left(v_{j}\,\frac{\partial}{\partial v_{j}}(A\,\psi_{A,b}(v))_{k}\right)_{1\leq j,k\leq d}\,=\,\frac{1}{2}\,A\,{\rm diag}(\beta_{1}v^{a_{1}}-\beta_{1}^{-1}v^{-a_{1}},\ldots,\beta_{n}v^{a_{n}}-\beta_{n}^{-1}v^{-a_{n}})\,A^{t}.$
This is the usual Jacobi matrix, with $\tfrac{\partial}{\partial v_{j}}$ replaced by the Euler operator $v_{j}\,\tfrac{\partial}{\partial v_{j}}$. One checks that for a point $v\in{\rm pr}_{\omega}^{-1}(\omega)$, the toric Jacobian determinant $\det J_{A,b}(v)$ vanishes if and only if the usual Jacobian determinant vanishes. We prefer the toric version because of the elegant expression $J_{A,b}(v)=\tfrac{1}{2}AD(v)A^{t}$, where $D(v)$ is the diagonal $n\times n$ matrix shown above.
Lemma 6.1.
The toric Jacobian $\det J_{A,b}(v)$ is not identically zero as a Laurent polynomial in $\beta$ and $v$. Moreover, for generic $\beta\in(\mathbb{C}^{*})^{n}$ and generic $\omega\in\mathbb{C}^{d}$, the fiber ${\rm pr}_{\omega}^{-1}(\omega)$ is finite. That is, for generic $\beta,\omega$ the equations (22) have finitely many solutions $v\in(\mathbb{C}^{*})^{d}$.
Proof 6.2.
For $v={\bf 1}\in(\mathbb{C}^{*})^{d}$ and $\beta=-i\cdot{\bf 1}\in(\mathbb{C}^{*})^{n}$, we have $\det J_{A,{\bf 1}}({\bf 1})=\det(-i\,A\,A^{t})\neq 0$ because $A$ has full rank. This shows the first statement. For the second part of the lemma, since $\beta$ is generic we may assume that $\det J_{A,b}(v)$ is not identically zero as a Laurent polynomial in $v$. Pick $v_{0}\in(\mathbb{C}^{*})^{d}$ such that $\det J_{A,b}(v_{0})\neq 0$ and let $\omega_{0}=A\psi_{A,b}(v_{0})$. By construction, $v_{0}$ is isolated in ${\rm pr}_{\omega}^{-1}(\omega_{0})$. Since ${\rm pr}_{\omega}$ is a dominant morphism of irreducible $d$-dimensional varieties, this implies that its generic fiber is finite [19, Chapter 1, §8, Theorem 2 and Corollary 1].
Remark 6.3.
For fixed $b$, the toric Jacobian $\det J_{A,b}$ might be identically zero. This happens for $A=\begin{pmatrix}1&1\end{pmatrix}$ when $\beta_{1}+\beta_{2}=0$. The equations (22) have no solutions for generic $\omega$.
Definition 6.4.
The ramification locus $R_{A,b}\subseteq W_{A,b}$ is the divisor
$R_{A,b}\,=\,\{(v,\omega)\in W_{A,b}\,:\,\det J_{A,b}(v)=0\}.$
The Lissajous discriminant $\nabla_{A,b}$ is the associated branch locus: $\nabla_{A,b}=\overline{{\rm pr}_{\omega}(R_{A,b})}\subseteq\mathbb{C}^{d}$. If $\nabla_{A,b}$ is a hypersurface, then its defining equation is denoted by $\Delta_{A,b}\in\mathbb{C}[\omega_{1},\ldots,\omega_{d}]$. If $\nabla_{A,b}$ has codimension greater than one, then we set $\Delta_{A,b}=1$.
Note that the polynomial $\Delta_{A,b}$ is defined up to scaling by a nonzero complex number.
Remark 6.5.
The Lissajous discriminant can be viewed as the branch locus of the linear projection of the Lissajous variety ${\cal L}_{A,b}$ given by the matrix $A$. It is an analog of the entropic discriminant [22], for which ${\cal L}_{A,b}$ is replaced by the reciprocal linear space of ${\rm Row}(A)$.
By definition, the Lissajous discriminant is the variety of the elimination ideal
$\langle\,A\,\psi_{A,b}(v)\,-\,\omega,\,\det(J_{A,b})\,\rangle\cap\mathbb{C}[\omega].$ (24)
Here we start from an ideal with $d+1$ generators in $\mathbb{C}[v^{\pm 1},\omega]$. We compute three examples.
Example 6.6.
For $A=\begin{pmatrix}1&2\end{pmatrix}\in\mathbb{Z}^{1\times 2}$, Equation˜24 reads
$\langle\beta v+\beta^{-1}v^{-1}+2(\beta^{2}v^{2}+\beta^{-2}v^{-2})-2\,\omega,\beta v-\beta^{-1}v^{-1}+4(\beta^{2}v^{2}-\beta^{-2}v^{-2})\rangle\cap\mathbb{C}[\omega].$
Setting $b=1$, $\beta=-i$, we compute $\Delta_{A,1}=256\omega^{4}-2367\omega^{2}+3375$, which has four real roots. A root $\omega^{(j)}$ of $\Delta_{A,1}$ corresponds to a tangent line $x+2y=\omega^{(j)}$ of ${\cal S}_{A}={\cal L}_{A,1}$; see Figure 6. For $b=0$, the discriminant is $\Delta_{A,0}=16\omega^{3}-31\omega^{2}-84\omega+99=(\omega-1)(\omega-3)(16\omega+33)$.
Figure 6: When $A=\begin{pmatrix}1&2\end{pmatrix}$, each of the four real roots $\omega^{(1)},\dotsc,\omega^{(4)}$ of $\Delta_{A,1}$ corresponds to a tangent line $x+2y=\omega^{(j)}$ (in red) of ${\cal S}_{A}$ (in blue).
Example 6.7.
As mentioned above, studying the Lissajous discriminant of the matrix $A$ is closely related to the bifurcation analysis of the dynamical system (20). Indeed, in view of linear stability analysis, an eigenvalue of the Jacobian matrix evaluated at a critical point can only change sign if its determinant vanishes. This happens when $\omega$ lies on the discriminant. For instance, the ramification locus $R_{A,b}$ in Example 5.7 is defined by the equations
$(v+v^{-1})-i(v-v^{-1})-2\,\omega\,=\,v-v^{-1}-i(v+v^{-1})\,=\,0,$
from which we see that $\nabla_{A,b}=\{\pm\sqrt{2}\}$. These discriminant points correspond to the two red lines $\{x+y=\pm\sqrt{2}\}$ tangent to the circle ${\cal L}_{A,b}(\mathbb{R})$, see Figure 5. The bounded discriminant chamber $(-\sqrt{2},\sqrt{2})$ contains ${\cal L}_{A,b}^{+}=(-1,1)$, and for $\omega$ in that chamber the dynamical system has one stable and one unstable equilibrium. For $\omega^{2}>2$, the system is unstable.
Example 6.8.
Let $A$ be as in Example 1.3. We compute the Lissajous discriminants of ${\cal L}_{A,b}$ for $b=\bf 0$ and $b=\bf 1$. The curve $\nabla_{A,\bf 0}$ has degree $6$, and $\nabla_{A,\bf 1}$ has degree $12$. The equations are
[…]
$+8\,\omega_{1}^{2}\,\omega_{2}^{3}-69\,\omega_{1}^{2}\,\omega_{2}^{2}+6\,\omega_{1}^{2}\,\omega_{2}+36\,\omega_{1}^{2}+20\,\omega_{1}\,\omega_{2}^{4}-46\,\omega_{1}\,\omega_{2}^{3}-6\,\omega_{1}\,\omega_{2}^{2}+36\,\omega_{1}\,\omega_{2}$
$+8\,\omega_{2}^{5}-23\,\omega_{2}^{4}-4\,\omega_{2}^{3}+36\,\omega_{2}^{2},$
[…]
where $e_{2}=\omega_{1}\omega_{2}+\omega_{1}\omega_{3}+\omega_{2}\omega_{3}$, $e_{3}=\omega_{1}\omega_{2}\omega_{3}$ are the elementary symmetric polynomials and $\omega_{3}=-\omega_{1}-\omega_{2}$. We will explain the symmetric structure of $\Delta_{A,\bf 1}$ in Proposition 6.14. Figure 7 shows our two discriminant curves. They are the branch loci of the projection of Figure 1 given by $A$. The reference [2, Section 5.3] studies $\nabla_{A,\bf 1}$ via machine learning techniques.
Different regions in Figure 7 correspond to different numbers of real solutions to (22), or different numbers of real intersection points in ${\cal L}_{A,b}\cap\{Ax=\omega\}$. For $\omega$ in the connected component of $\mathbb{R}^{2}\setminus\nabla_{A,\bf 0}$ highlighted in orange, the line $Ax=\omega$ intersects ${\cal C}_{A}$ in three real points. In all other connected components, there is only one real intersection point. On the right, the green, blue and red components correspond to $\omega$ with six, four and two real intersection points respectively. We note that the convex hull of the real points of $\nabla_{A,\bf 1}$ contains the blue region $\Omega^{+}$ in Figure 3 (right), but the two do not coincide. Figure 8 shows a comparison. This is similar to the inclusion $(-1,1)\subsetneq(-\sqrt{2},\sqrt{2})$ in Example 6.7.
Figure 7: Lissajous discriminants from Example 6.6.
Figure 8: Comparing $\Omega^{+}$ (blue) with the Lissajous discriminant (purple).
Theorem 6.9.
If the Lissajous discriminant variety $\nabla_{A,b}$ is a hypersurface, then it has degree at most $(\deg{\rm pr}_{\omega})^{-1}\cdot d\cdot d!\,{\rm vol}(P_{A})$, where ${\rm pr}_{\omega}$ denotes the restriction $({\rm pr}_{\omega})_{|R_{A,b}}$. The number ${\rm vol}(P_{A})$ is as in Theorem 2.5.
Proof 6.10.
For generic $\omega_{0},\omega_{1}\in\mathbb{C}^{d}$, the degree of $\nabla_{A,b}$ is the number of intersection points of $\nabla_{A,b}$ with the line parametrized by $\omega_{0}+t\,\omega_{1}$. This is at most the number of solutions $(v,t)\in(\mathbb{C}^{*})^{d+1}$ to the system of $d+1$ Laurent polynomial equations
$A\,\psi_{A,b}(v)\,=\,\omega_{0}+t\,\omega_{1},\quad\det J_{A,b}(v)\,=\,0.$ (25)
Let us denote this number by $\delta$. In fact, it is clear that if $(v,t)$ is a solution to these equations, then so is $(v^{\prime},t)$ for any $v^{\prime}\in{\rm pr}_{\omega}^{-1}(\omega_{0}+t\,\omega_{1})$. Hence, we have $\deg(\nabla_{A,b})=(\deg{\rm pr}_{\omega})^{-1}\cdot\delta$.
By Bernstein’s theorem, the number $\delta$ is bounded by the mixed volume of the $d+1$ Newton polytopes of the equations [3]. The Newton polytope of each of the first $d$ equations is contained in the $(d+1)$-dimensional polytope $\hat{P}_{A}={\rm conv}(P_{A}\cup e_{d+1})$. Here we add the $(d+1)$-st standard basis vector to $P_{A}$ because of the term $t\,\omega_{1}$. The Newton polytope of the toric Jacobian determinant $\det J_{A,b}(v)$ is contained in the $d$-dimensional polytope $d\cdot P_{A}$. Hence, we have $\delta\leq{\rm MV}(\hat{P}_{A},\ldots,\hat{P}_{A},d\cdot P_{A})$, where $\hat{P}_{A}$ is repeated $d$ times and ${\rm MV}(\cdot)$ denotes the mixed volume. By definition, this mixed volume equals the coefficient of $\lambda_{0}^{d}\lambda_{1}$ in ${\rm vol}(\lambda_{0}\hat{P}_{A}+\lambda_{1}dP_{A})$. The Minkowski sum $\lambda_{0}\hat{P}_{A}+\lambda_{1}dP_{A}$ is obtained from the pyramid $(\lambda_{0}+d\lambda_{1})\hat{P}_{A}$ with volume $(\lambda_{0}+d\lambda_{1})^{d+1}(d+1)^{-1}{\rm vol}(P_{A})$ by “chopping off” the top of the pyramid $\lambda_{0}e_{d+1}+d\lambda_{1}\hat{P}_{A}$ with volume $(d\lambda_{1})^{d+1}(d+1)^{-1}{\rm vol}(P_{A})$. The coefficient standing with $\lambda_{0}^{d}\lambda_{1}$ in the difference of these expressions is $d\cdot d!{\rm vol}(P_{A})$. We have shown that $\delta\leq d\cdot d!{\rm vol}(P_{A})$, and since $\deg(\nabla_{A,b})=(\deg{\rm pr}_{\omega})^{-1}\cdot\delta$ the theorem follows.
Example 6.11.
The degree of the Lissajous discriminant can be computed by solving (25) for generic $\omega_{0},\omega_{1}$. There are $\delta$ solutions, whose image under ${\rm pr}_{\omega}$ consists of $\deg\nabla_{A,b}$ many points. To compute all $\delta$ solutions, we use the package HomotopyContinuation.jl (v2.15.0) [4] in Julia (v1.11). To compute the lattice volume $d!\,{\rm vol}(P_{A})$, we use the function lattice_volume from Oscar.jl (v1.1.2) [8, 20]. We apply these methods for matrices $A$ coming from cyclic and complete graphs. The results are reported in Table 1. The column $d\cdot d!\,{\rm vol}(P_{A})$ is computed using the command d*lattice_volume(convex_hull(transpose([A -A]))). By [5, Theorem 14], these numbers for $C_{n}$ are given by the sequence $n(n-1)\binom{n-1}{\lfloor\tfrac{1}{2}(n-1)\rfloor}$. The other entries can be reproduced using the following Julia code snippet, which uses standard tools:
using HomotopyContinuation
A = ... # input a matrix A of full row rank, e.g., A = [1 0 -1; -1 1 0]
b = ... # input a vector b, e.g., b = [0;0;0]
d, n = size(A)
β = exp.(-im*b*pi/2); βinv = β.^(-1);
@var v[1:d] w[1:d] t # declare variables
Ad = [[maximum([aa,0]) for $a$ in A];-[minimum([aa,0]) for $a$ in A]]
y = [prod([v;w].^Ad[:,i]) for i = 1:n]
yinv = [prod([w;v].^Ad[:,i]) for i = 1:n]
ψ = [1//2*(β[i]*y[i] + βinv[i]*yinv[i]) for i = 1:n]
D = det(A*diagm([1//2*(β[i]*y[i]-βinv[i]*yinv[i]) for i = 1:n])*transpose(A))
myω0 = randn(ComplexF64,d); myω1 = randn(ComplexF64,d);
eqs = System([A - myω0+t*myω1; [v[i]*w[i]-1 for i = 1:d]; D], variables = [v;w;t])
R = HomotopyContinuation.solve(eqs)
δ = length(solutions(R))
degdisc = length(unique_points([[sol[end]] for sol in solutions(R)]))
degpr = δ/degdisc
The variables in line 6 are our variables $v_{1},\ldots,v_{d}$ and the variable $t$ in Equation 25. The variables $w_{k}$ play the role of the inverses of the $v_{k}$, as in Section 2. The columns of the matrix Ad in line 7 are the concatenations of nonnegative integer vectors $a_{j}^{+}$ and $a_{j}^{-}$ satisfying $a_{j}=a_{j}^{+}-a_{j}^{-}$. Lines 8-13 construct the system of equations (25), and line 14 solves it. We emphasize that numerical homotopy continuation methods are based on heuristics and do not provide a proof that the numbers in the table are correct. However, we expect they are. Our goal here is to exemplify the (non-)tightness of the bound in Theorem 6.9.
AA $d\cdot d!{\rm vol}(P_{A})$ $b$ generic $b=\mathbf{0}$ $b=\mathbf{1}$
$\deg(\nabla_{A,b})$ $\deg(\mathrm{pr}_{\omega})$ $\deg(\nabla_{A,\bf 0})$ $\deg(\mathrm{pr}_{\omega})$ $\deg(\nabla_{A,\bf 1})$ $\deg(\mathrm{pr}_{\omega})$
$A(C_{3})$ $12$ $12$ 11 66 22 $12$ 11
$A(C_{4})$ $36$ $36$ 11 $\bf 12$ 22 $\bf 12$ 22
$A(C_{5})$ $120$ $120$ 11 $60$ 22 $120$ 11
$A(C_{6})$ $300$ $300$ 11 $\bf 130$ 22 $150$ 22
$A(C_{7})$ $840$ $840$ 11 $420$ 22 $840$ 11
$A(C_{8})$ $1960$ $1960$ 11 $\bf 910$ 22 $\bf 910$ 22
$A(K_{4})$ $60$ $60$ 11 $\bf 26$ 22 $\bf 48$ 11
$A(K_{5})$ $280$ $280$ 11 $\bf 90$ 22 $\bf 140$ 11
$A(K_{6})$ $1260$ $1260$ 11 $\bf 276$ 22 $\bf 360$ 11
Table 1: $\deg(\nabla_{A,b})$ for different reduced incidence matrices $A=A(G)$. Numbers in bold indicate that the upper bound from Theorem 6.9 is not attained.
Theorem 6.9 and Example 6.11 show that Lissajous discriminants may have large degrees, which makes them challenging to compute. In some cases, the polynomial $\Delta_{A,b}$ exhibits some symmetries, and this can be exploited in the computation. Here is an example.
Proposition 6.12.
The discriminant $\Delta_{A,\bf 1}$ satisfies $\Delta_{A,\bf 1}(\omega)=\alpha\Delta_{A,\bf 1}(-\omega)$, where $\alpha=\pm 1$. In particular, the degree of each monomial appearing in $\Delta_{A,\bf 1}$ is even if $\alpha=1$, or odd if $\alpha=-1$.
Proof 6.13.
The Lissajous discriminant is stable under changing the sign of $\omega$. Indeed, we have
[…]
This implies that $\Delta_{A,\bf 1}(\omega)=\alpha\,\Delta_{A,\bf 1}(-\omega)$ for some $\alpha\in\mathbb{C}^{*}$. Changing signs twice, we find $\Delta_{A,\bf 1}(\omega)=\alpha\,\Delta_{A,\bf 1}(-\omega)=\alpha^{2}\,\Delta_{A,\bf 1}(\omega)$, hence $\alpha=\pm 1$. Finally, write $\Delta_{A,\bf 1}=\Delta_{e}+\Delta_{o}$, where $\Delta_{e}$ is the sum of the monomials of $\Delta_{A,\bf 1}$ with even degree, and $\Delta_{o}$ contains those with odd degree. If $\alpha=1$, then $2\Delta_{o}=\Delta_{A,\bf 1}(\omega)-\Delta_{A,\bf 1}(-\omega)=0$. Similarly, if $\alpha=-1$, then $\Delta_{e}=0$.
If $A$ comes from a graph $G$, as in Section 4, then the Lissajous discriminant respects the symmetries of $G$. We shall now make this more precise. Recall that an automorphism of $G$ is a permutation of its vertices which preserves adjacency. These constitute the automorphism group ${\rm Aut}(G)$, a subgroup of the symmetric group of order $n!$. For instance, the automorphism group of the complete graph $K_{n}$ is ${\rm Aut}(K_{n})=S_{n}$, the full symmetric group. The automorphism group of the cycle graph $C_{n}$ is the dihedral group of order $2n$.
Previously, we have worked with the reduced incidence matrix $A(G)\in\mathbb{Z}^{d\times n}$, which has full rank $d$. To describe the symmetries of the discriminant, it is more natural to work with the full incidence matrix $A_{G}\in\mathbb{Z}^{(d+1)\times n}$ of rank $d$ and consider the equivalent equations
$A_{G}\,\psi_{A_{G},\bf 1}(v)=\omega,\,\,v_{d+1}=1\quad\text{and}\quad{\rm rank}(J_{A_{G},b}(v))<d.$
Since the rows of $A_{G}$ sum to zero, these equations imply that $\omega_{1}+\cdots+\omega_{d}+\omega_{d+1}=0$. The Lissajous discriminant $\nabla_{A_{G},\bf 1}$ is (expected to be) a hypersurface inside the hyperplane $\omega_{1}+\cdots+\omega_{d}+\omega_{d+1}=0$. Its projection onto the first $d$ coordinates is $\nabla_{A(G),\bf 1}$.
Proposition 6.14.
Let $G$ be a graph with $m=d+1$ vertices and let $\sigma\in{\rm Aut}(G)$. We have $\omega\in\nabla_{A_{G},\bf 1}\subset\mathbb{C}^{d+1}$ if and only if $\sigma(\omega)\in\nabla_{A_{G},\bf 1}$, where $\sigma$ acts on $\omega$ by permuting coordinates.
Proof 6.15.
The group ${\rm Aut}(G)$ acts on $(\mathbb{C}^{*})^{d+1}$ and on $\mathbb{C}^{d+1}$ by permuting coordinates. The proposition follows from the observation that the map $f_{G}:(\mathbb{C}^{*})^{d+1}\rightarrow\mathbb{C}^{d+1}$ given by $v\mapsto A_{G}\,\psi_{A_{G},\bf 1}(v)$ is equivariant with respect to these actions, meaning that $f_{G}(\sigma(v))=\sigma(f_{G}(v))$, $\sigma\in{\rm Aut}(G)$. To prove this, write the $k$-th coordinate of $f_{G}(v)$ as
$f_{G}(v)_{k}\,=\,\sum_{(k,j)>0}\tfrac{1}{2i}(v_{k}v_{j}^{-1}-v_{k}^{-1}v_{j})-\sum_{(k,j)<0}\tfrac{1}{2i}(v_{j}v_{k}^{-1}-v_{j}^{-1}v_{k}),$
where $(k,j)>0$ denotes a sum over all edges $(k,j)$ oriented from $k$ to $j$, and similarly for $(k,j)<0$. This simplifies to $f_{G}(v)_{k}=\sum_{(k,j)\in E_{k}}\tfrac{1}{2i}(v_{k}v_{j}^{-1}-v_{k}^{-1}v_{j})$, where $E_{k}$ is the set of all edges adjacent to vertex $k$. Now the equality $f_{G}(\sigma(v))_{k}=f_{G}(v)_{\sigma(k)}$ is clear from
$\sum_{(k,j)\in E_{k}}\tfrac{1}{2i}(v_{\sigma(k)}{v_{\sigma(j)}}^{-1}-{v_{\sigma(k)}}^{-1}v_{\sigma(j)})=\sum_{(\sigma(k),\sigma(j))\in E_{\sigma(k)}}\tfrac{1}{2i}(v_{\sigma(k)}{v_{\sigma(j)}}^{-1}-{v_{\sigma(k)}}^{-1}v_{\sigma(j)}),$
since $(k,j)\in E_{k}$ if and only if $(\sigma(k),\sigma(j))\in E_{\sigma(k)}$, by definition of ${\rm Aut}(G)$.
Example 6.16.
Expanding $\Delta_{A({C_{3}}),\bf 1}$ from Example 6.6 as a polynomial in $\omega_{1}$ and $\omega_{2}$, we obtain 41 terms, all of even degree. The fact that $\Delta_{A_{C_{3}},\bf 1}(\omega_{1},\omega_{2},\omega_{3})$ can be expressed in terms of elementary symmetric polynomials mirrors the fact that ${\rm Aut}(C_{3})=S_{3}$.